### v1

# Probing Visual Language Priors in VLMs

Abstract
Despite recent advances in Vision-Language Models (VLMs), many still over-rely on visual language priors existed in their training data rather than true visual reasoning. To examine the situation, we introduce ViLP, a visual question answering (VQA) benchmark that pairs each question with three potential answers and three corresponding images: one image whose answer can be inferred from text alone, and two images that demand visual reasoning. By leveraging image generative models, we ensure significant variation in texture, shape, conceptual combinations, hallucinated elements, and proverb-based contexts, making our benchmark images distinctly out-of-distribution. While humans achieve near-perfect accuracy, modern VLMs falter; for instance, GPT-4 achieves only 66.17% on ViLP. To alleviate this, we propose a self-improving framework in which models generate new VQA pairs and images, then apply pixel-level and semantic corruptions to form “good-bad” image pairs for self-training. Our training objectives compel VLMs to focus more on actual visual inputs and have demonstrated the effectiveness in enhancing the performance of open-source VLMs, including LLaVA-v1.5 and Cambrian.
1 Introduction
Vision-Language Models (VLMs) have advanced text-image interaction, serving as a cornerstone in bridging the gap between visual and textual data (Achiam et al., 2023; Team et al., 2023). However, a persistent challenge for learning-based models, such as ResNets and CLIPs, is their reliance on strong priors overly learned in the training data, sometimes overlooking visual cues for answering questions (Agrawal et al., 2016; Prabhu et al., 2023). For example, when shown an image of a hexagonal moon (Figure 1), a model might incorrectly assert that “the moon is round" due to strong language priors. Similarly, these models may adhere to learned visual priors (Thrush et al., 2022; Sterz et al., 2024), making it difficult for them to comprehend out-of-distribution visual cues, such as a zebra with atypical spot patterns (Figure 1), where humans would easily discern the difference. This raises an important question: do today’s VLMs still over-rely on learned visual language priors?
To investigate this, we probe the visual language priors of VLMs by constructing Question-Image-Answer (QIA) triplets that deliberately deviate from the training data distribution. Unlike existing benchmarks that typically rely on internet-sourced images, which inadvertently favor the visual language priors embedded in training data, we utilize modern image generation models such as DALL·E-3 (Ramesh et al., 2021) and Flux to synthesize out-of-distribution images, supporting uncommon answers. To further challenge the priors of VLMs, we amplify the language priors in questions by introducing distractor facts, making the questions more suggestive.
Our benchmark, ViLP, contains 300 carefully designed questions, each paired with three distinct answers: a Prior Answer and two Test Answers, resulting in a total of 900 QIA triplets. Each question is structured to present a distractor fact followed by a question. The distractor fact directly leads to the Prior Answer. In contrast, the two Test Answers are crafted to challenge the priors by requiring both textual and visual cues for accurate reasoning. While human participants can differentiate between all three answers, current VLMs face considerable difficulty, as evidenced by a significant performance drop in our benchmarks, with GPT-4o (OpenAI, 2024) scoring only 66.17%.
Inspired by ViLP  results, we propose a self-improving approach to encourage VLM reasoning more relying on visual inputs. Our method involves self-generating VQAs using image generation and editing models (Podell et al., 2023; Ren et al., 2024; Brooks et al., 2023a) and introducing controlled corruptions to create “good" and “bad" image pairs for DPO-like training (Rafailov et al., 2024). We theoretically prove that our proposed objective optimizes an upper bound of the RLHF objective (Ouyang et al., 2022), and experiments on open-sourced VLMs, including LLaVA-v1.5 (Liu et al., 2024) and Cambrian (Tong et al., 2024), demonstrate its effectiveness.
Figure 1: Sample data from ViLP. For the same question, ViLP  provides three distinct images and corresponding answers (upper-left corner). All questions follow a consistent structure, combining a distractor fact with a question. The Prior Answer (first column) can be directly inferred from the question, while Test Answers (second &\&& third column) rely on visual cues. All of our answers are designed to be single words, allowing for evaluation without the need for LLMs. We have developed a robust synonym and plural detection pipeline to ensure that open-ended responses do not interfere with the evaluation process. Please refer to Appendix A for more data samples from ViLP.
2 Related Work
VQA Dataset: There have been long-term efforts to establish Visual Question Answering (VQA) datasets from various perspectives, including general visual question answering  (Agrawal et al., 2015; Gurari et al., 2018; Fu et al., 2023; Liu et al., 2023e; Li et al., 2023a; Yu et al., 2023b; Liu et al., 2024), reading text or charts (Singh et al., 2019a; Mathew et al., 2020; 2021; Masry et al., 2022), complicated reasoning (Lu et al., 2022; 2023), probing compositions (Hudson & Manning, 2019; Ma et al., 2022; Thrush et al., 2022; Hsieh et al., 2023; Li et al., 2024a), studying hallucinations (Rohrbach et al., 2018; Li et al., 2023c; Guan et al., 2023), common sense reasoning (Bitton-Guetta et al., 2023b; a), and more (Majumdar et al., 2024; Sterz et al., 2024). This paper proposed a benchmark to evaluate the visual reasoning abilities of VLMs when the questions, answers and images defy common visual language priors. Our approach follows the balanced dataset design in Goyal et al. (2017), where each question is associated with three answers: the first aligns with language priors, while the latter two deviate from them, requiring reliance on visual cues for correct answering. By leveraging state-of-the-art image generation models, our benchmark creates images that defy priors much more effectively than previous works using internet images, such as Goyal et al. (2017); Tong et al. (2024). Moreover, unlike the “trick” category in Sterz et al. (2024), our process first generates question-answer pairs and then synthesizes images that meet the specified conditions, resulting in more challenging visual reasoning tasks. More comparisons with existing datasets are included in Appendix D.2.
Vision Language Models and Language Priors: Multimodal reasoning is essential for machine intelligence, with vision-language models (VLMs) integrating visual perception, text reasoning, instruction following, and generation for complex tasks (Tan & Bansal, 2019; Li et al., 2019; Kim et al., 2021; Wang et al., 2021b; a; Alayrac et al., 2022; Li et al., 2023b; Chen et al., 2022; Jia et al., 2021; Shen et al., 2021; Singh et al., 2021; Liu et al., 2023c; a; Zhao et al., 2023; Chen et al., 2023; Zhu et al., 2024; Li et al., 2024c; Dai et al., 2023; Li et al., 2024c; Yu et al., 2024; Dai et al., 2024; Deitke et al., 2024). Building on the success of language models (Brown et al., 2020; OpenAI, 2023a; Touvron et al., 2023a; b; Chiang et al., 2023) and pre-trained visual encoders (Radford et al., 2021; Desai & Johnson, 2020; Caron et al., 2021; Chen et al., 2024), many recent advancements leverage vision-language paired datasets (Liu et al., 2024; Tong et al., 2024) to fine-tune intermediate connectors between LLMs and visual encoders (Liu et al., 2024). However, these paired datasets are significantly smaller than the large text corpora used for LLM pre-training (OpenAI, 2023b; Soldaini et al., 2024), and many approaches do not update the visual encoder and LLM parameters during fine-tuning, which can lead to persistent language biases. Specifically, the outputs of VLMs become dominated by the language priors learned from the pre-training corpus, reducing the importance of visual inputs (Thrush et al., 2022; Sterz et al., 2024). This issue becomes even more critical when VLMs are presented with deliberately generated images that challenge these biases, as demonstrated in our work. Previous methods and datasets (Goyal et al., 2016; Agrawal et al., 2017; Dancette et al., 2021; Wu et al., 2022; Ramakrishnan et al., 2018; Gouthaman & Mittal, 2020) have attempted to assess these issues by studying visual backbones and curating datasets using simulators (Johnson et al., 2016) or internet images (Zhang et al., 2016). This paper examines these challenges by creating a novel VQA benchmark, featuring carefully designed questions, fact-based distractors, and rare-distribution answers. Furthermore, we utilize advanced image generation techniques to produce realistic images that subvert common language biases as shown in Figure 1.
Self-Rewarding VLM: Self-rewarding LLM (Yuan et al., 2024) has demonstrated the potential of using LLMs to generate and optimize data through directed reference optimization (Rafailov et al., 2024), leading to self-improvements. This line of research has been extended to VLMs, where new answers are generated and rated for DPO training (Zhou et al., 2024a; Deng et al., 2024; Zhou et al., 2024c; Wang et al., 2024b; a). Our findings resonate with this body of work on self-rewarding VLMs; however, we present two key distinctions: (1) our proposed image-DPO method generates multiple images for a single question-image pair, rather than generating multiple answers and associated ratings; (2) unlike previous methods that focused solely on existing images, both our proposed image-DPO and text-DPO pipelines produce a diverse array of new images with pre-trained image generative models, including SDXL (Podell et al., 2023), GroundedSAM (Ren et al., 2024), and InstructPix2Pix (Brooks et al., 2023a). Specifically, as illustrated in Figure 6, our proposed text-DPO is similar to the self-rewarding VLMs mentioned earlier but incorporates the generation of new images using pre-trained image generation models. In contrast, our image-DPO deliberately corrupts images to create multiple degraded versions, which are used as rejected data during DPO training. This is intended to compel the VLM to rely more on the visual input for accurate generation.
3 ViLP  Benchmark
3.1 Design Principles
"What’s the tall animal with a long neck?" Even without any visual cues, humans readily infer that the answer is "giraffe", leveraging language context and commonsense knowledge. But is it necessarily a giraffe? As shown in the bottom-right of Figure 1, there may be an elephant or a camel with a very long neck and much taller than giraffes. This example highlights a potential issue with current large Vision-Language Models (VLMs). As VLMs are built upon Large Language Models (LLMs) trained on web-scale corpora and are fine-tuned with comparatively much smaller image-text datasets, they may develop a strong bias toward certain visual language priors (especially language) presented in training data and overlook real visual input (Thrush et al., 2022; Sterz et al., 2024). Regarding visual language priors, we focus on strong language priors that confuse VLMs to derive direct answers solely from textual cues. Simultaneously, we aim to create images that are distinctly out-of-distribution to challenge potential visual priors which make the model overlook the critical visual cues such as the unusual spots patterns in the Zebra shown in Figure 1.
To assess the visual reasoning abilities of VLMs against visual language priors, we introduce a new benchmark, ViLP, featuring carefully constructed out-of-distribution question-image-answer (QIA) triplets. These triplets are designed based on the below principles ensuring that VLMs rely on both textual and visual cues, rather than learned priors, to arrive at correct answers.
  • Text-Only Inference: Create questions that, when considered alone, lead to a high-confidence answer based solely on textual information.
  • Visual Influence: When an accompanying image is present, the correct answer differs significantly from the text-only response, sometimes contradicting common sense, ensuring the critical role of visual context in ensuring a high-confidence result.
Mathematically, with $Q$ denoting a question, $I$ denoting an image, $A=\{a_{\text{prior}},\cdots,a_{\text{test}},\cdots\}$ being the set of all possible answers, we use $P(a\mid Q)$ and $P(a\mid Q,I)$ to denote the probability of answer $a$ given the question $Q$ alone, and the probability of answer $a$ given both the question $Q$ and the image $I$, respectively. Here $p$ is a prior model, which could either be human’s cognition prior ($P_{\text{human}}$) or a VLM/LLM’s learned visual language prior ($P_{\theta}$). For constructing our benchmark, we used the following objectives:
Criteria One: Without considering the image, the question $Q$ should lead to a high-confidence answer $a_{\text{prior}}$, where $\delta_{1}$ is a high-confidence threshold. $a_{\text{prior}}$ usually satisfies common knowledge, such as “the moon is round" and “Leonardo da Vinci painted the Mona Lisa" as shown in Figure 1.
$P(a_{\text{prior}}\mid Q)\geq\delta_{1}$ (1)
Criteria Two: With the inclusion of the image $I$, the correct answer becomes $a_{\text{test}}$, where $\delta_{2}$ is another high-confidence threshold. Also, $a_{\text{test}}$ is significantly different from $a_{\text{prior}}$ to ensure that the image has a substantial impact on the answer, where $D$ is a divergence measure (e.g., Kullback-Leibler divergence), and $\delta_{3}$ is a threshold indicating significant difference. For instance, the image in the second-to-last row and column of Figure 1 clearly shows the moon as a hexagon.
$P(a_{\text{test}}\mid Q,I)\geq\delta_{2}\quad\text{and}\quad D\big{(}P(a_{\text{prior}}\mid Q),P(a_{\text{test}}\mid Q,I)\big{)}\geq\delta_{3}$ (2)
Criteria Three: We design answer $a_{\text{test}}$ to be a rare choice, one that is unlikely to be inferred from the question $Q$ alone. In contrast, answer $a_{\text{prior}}$ should be clearly incorrect when considering both the question $Q$ and the image $I$, where $\delta_{4}$ represents a low-confidence threshold. Typically, $a_{\text{test}}$ will be out of context. For example, in the image from the 3rd row and 2nd column of Figure 1, Vincent van Gogh serves as $a_{\text{test}}$, inferred from the image, while Leonardo becomes contradictory after incorporating the additional image condition.
$P(a_{\text{test}}\mid Q)\leq\delta_{4}\quad\text{and}\quad P(a_{\text{prior}}\mid Q,I)\leq\delta_{4}$ (3)
In constructing ViLP, we use human cognition prior $P_{\text{human}}$ as guidance for its design. We evaluated the behavior of the learned priors of VLMs/LLMs by benchmarking $P_{\theta}$ against the human prior $P_{\text{human}}$. This benchmarking process implicitly measuring the divergence between $P_{\text{human}}$ and $P_{\theta}$.
3.2 Question-Image-Answer Generation
As outlined in Criteria Three, we aim for the test answer, $a_{\text{test}}$, to appear highly improbable when considered solely in the context of the question $Q$, yet become a clear and confident choice when paired with the image $I$. This objective presents a significant challenge in finding images $I$, as such images may not follow common knowledge and not naturally exist in the real world or the internet. Recent advancements in generative models offer a solution. By utilizing advanced image generation models such as DALL·E-3 (Ramesh et al., 2021) and Flux, which follow text prompts to synthesize images, we can create images that blend unusual elements. This capability allows us to generate images with the necessary visual coherence to satisfy the specific demands of the question and answer, where we create in a way to disobey language priors. To further ensure the generation of question-image-answers (QIAs) that align with the principles outlined in Section 3.1, our benchmark dataset design incorporates substantial human effort, assisted by state-of-the-art language models such as OpenAI-o1 and Claude-3.5-Sonnet, with additional details provided in Appendix D.1.
Briefly, for each question, we designed three different answers where the first answer $a_{\text{prior}}$ could be inferred solely from the question text. The other two answers were specifically designed against language priors, challenging the model to interpret visual cues despite strong textual priors. During evaluation, we focus on the model’s performance on these second and third answers, i.e., $a_{\text{test}}$. After crafting the initial questions and answers, we leverage GPT-4 to generate text prompts that guide the creation of diverse images at scale. This facilitates the production of varied visual data for the next stage, where high-quality QIA triplets are selected through human review and filtering. Both the text and images are refined as necessary when misalignment or ambiguity are identified. We ensure that the final QIA triplets are clear and easily interpretable by humans, as evidenced by the results of our human evaluation results in Table 2. This iterative process involves several cycles of adjustments to achieve the desired quality. Throughout the process, we encountered two main challenges: (1) generating diverse, out-of-distribution, while reasonable question-answer pairs are difficult, as models like OpenAI-o1 and Claude-3.5-Sonnet tend to converge on a limited set of responses, requiring significant human input to introduce variety; and (2) generating images that against with specific priors using state-of-the-art image generation models is also challenging. In some cases, it was necessary to generate hundreds of images to find just one that accurately matched the question and answer, while avoiding ambiguity.
Table 1: Categorical information.
Type Frequency
Texture 16
Shape 20
Conceptual combinations 276
Hallucinated Components 151
Proverbs 17
As a result, we created a total of 300 questions, each paired with three distinct image-answer combinations, resulting in 900 IQA sets. These questions span a broad range of topics, from low-level visual recognition tasks such as texture and shape, to higher-level visual reasoning involving conceptual combinations, hallucinated components, and proverbs. To reinforce the text priors, we also provide a distractor fact before the question, offering additional context for the prior answer $a_{\text{prior}}$ (the one inferred without the image). Some categorical information is shown on the right Table 1. Most of our data are not confined to a single category; instead, they frequently combine multiple categories to challenge even the most advanced VLMs. For instance, in the moon example (6th-row, 2nd-col), we explore shapes and conceptual combinations, such as a round moon paired with a hexagon versus a hexagonal moon. Similarly, the umbrella example (7th-row, 1st-col) incorporates both conceptual combinations and hallucinated components, blending elements like a plunger and a blender. On average, our data span 1.6 categories per question.
3.3 Dataset Evaluation
All of our questions are designed to be answered with a single word, as this design greatly simplifies evaluation protocol compared to assessing full sentences, without requiring LLMs for evaluation. We explicitly instruct the model to provide a single-word answer, and we evaluate the correctness of each response using a binary system. To ensure a fair evaluation, we devote significant effort to building a comprehensive set of synonyms and plural for each answer to detect other valid alternative answers. This ensures that the model is only penalized for actual errors, not for providing synonymous or alternative correct responses.
4 Image DPO
Inspired by our benchmark results, we propose Image DPO, a self-improvement method for enhancing VLMs’ visual reasoning, featuring a new objective and a data generation pipeline using VLMs themselves and pre-trained image models.
4.1 Objective
Figure 2: Illustration of Image DPO. We construct chosen and rejected pairs by corrupting the image with a set of perturbations while keeping the text (questions and answers) unchanged. This setup encourages that the model more focus on the image to differentiate between the chosen and rejected pairs, as both pairs contain identical textual information. The perturbations applied to the images include semantic editing, Gaussian blurring, and pixelization. More details in Appendix C.2.
Existing approaches for VLM self-improvement usually follow the Direct Preference Optimization (DPO) (Rafailov et al., 2024) used in Large Language Models (LLMs), where the model is trained to distinguish between good and bad answers for a fixed image and question. However, this straightforward adaptation is not vision-centered, as the model can still distinguish good and bad answers from the text alone without needing to analyze the image. In contrast, we propose Image DPO, a vision-focused objective that creates good and bad question-image-answer pairs by corrupting the image while keeping the question and answer unchanged (Figure 2).
Mathematically, given the image $I_{w}$, question $Q$ and the corresponding answer $A$, we generate a corrupted image $I_{l}$ by various image editing operations including gaussian blur, pixelation or semantic editing. With the corrupted image $I_{l}$, the triplet of $Q$, $I_{l}$, and $A$ should form a bad question-image-answer pair compared to the original triplet of $Q$, $I_{w}$, and $A$. We train our model to distinguish between good and bad question-image-answer triplets using the objective function:
$L(\theta,\theta_{\text{ref}})=-\mathbb{E}_{Q,I_{w},I_{w},A\sim S}\left[\log\sigma(\alpha\frac{\pi_{\theta}(A\mid Q,I_{w})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{w})}-\alpha\frac{\pi_{\theta}(A\mid Q,I_{l})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{l})})\right]$ (4)
Here, $\theta$ represents the VLM model being trained, $\theta_{\text{ref}}$ is the reference VLM model (typically the previous version of $\theta$), $S$ is the training data consisting of good and bad question-image-answer triplets, $\sigma$ is the sigmoid function, and $\alpha$ is a positive constant.
In Appendix B, we demonstrate that this objective optimizes an upper bound of the RLHF objective (Ouyang et al., 2022), sets it apart from the original DPO objective. Intuitively, since the textual inputs and outputs are identical in both good and bad cases, the gradients of this objective push the model to rely more on the vision branch, driving a shift in gradient direction when processing high-quality images $I_{w}$ compared to corrupted images $I_{l}$, as illustrated in Appendix Figure 9. This behavior encourages the model to focus more on image inputs rather than relying solely on text-based reasoning, thereby enhancing its performance on visually-related tasks.
4.2 Data Generation
Training VLMs requires a large amount of question-image-answer (QIA) triplets, which are often scarce and expensive to collect. In response, we present a scalable data generation pipeline that creates new QIAs from existing datasets by leveraging VLMs and pre-trained image generation and editing models, as illustrated in Appendix Figure 10. Our pipeline is akin to the tool usage of VLMs/LLMs as prior works (Gupta & Kembhavi, 2022; Parisi et al., 2022; Yao et al., 2022; Sur’is et al., 2023), where pre-trained image models are invoked as callable functions. Given a seed image, VLMs are tasked with simultaneously selecting appropriate functions (e.g., image generation or editing models) and generating corresponding instructions. These instructions are then used to produce new images, in addition to the seed image, as illustrated in Figure 11. The same VLMs are then employed to generate QA pairs for these newly created images. Following, we apply the mentioned three types of image corruptions to the generated images, constructing good-bad QIA pairs (i.e., $I_{w}$ and $I_{l}$). Specifically, we employ Stable Diffusion XL (Podell et al., 2023; Rombach et al., 2022) for image generation, and use Instruct-Pix2Pix (Brooks et al., 2023a), and Grounded-SAM (Rombach et al., 2022; Ren et al., 2024) for image editing. All of our seed images are sourced from existing datasets, including COCO (Lin et al., 2014), Text2VQA (Singh et al., 2019b), and Visual Genome (Krishna et al., 2017). Comprehensive details are provided in Appendix C.
5 Experiments
Figure 3: Qualitative examples. We show the results from GPT-4o, Claude-3.5-Sonnet, Gemini-1.5-Pro, and Llama-3.2-Vision-90B for some challenging cases.
We introduce ViLP, a benchmark designed to evaluate VLMs’ visual reasoning capabilities in the presence of strong visual language priors, consisting of 300 carefully crafted questions. Each question is paired with three distinct images and corresponding answers, forming one QIAprior (QIA: question-image-answer) and two QIAtest, resulting in a total of 900 unique QIAs. The QIAprior aligns with common knowledge and language priors, making it easy to infer even from text alone. The QIAtest, on the other hand, challenges the VLM’s visual reasoning ability against these priors.
ViLP  includes two evaluation settings: ViLPF (with distractor facts provided) and ViLPP (pure questions without distractor facts). For each setting, we report the average accuracy on QIAtest as our focused benchmark Score, and the accuracy on QIAprior as Prior. Results are on the Table 2.
Is the QIA easy for humans? We first evaluate our benchmark using human study. Humans achieved nearly 100% accuracy on QIAprior and over 98% accuracy on QIAtest, demonstrating that our question-image-answer combinations are non-ambiguous for humans to answer. Despite QIAtest being designed as out-of-distribution examples, humans were able to distinguish them.
Humans performed slightly better on ViLPF-QIAprior when distractor facts were provided, as they could easily identify that these facts aligned with the correct answers. Moreover, Score on ViLPF-QIAtest was marginally lower when facts were introduced, as the distractor facts added some noise and caused minor confusion, although the impact of this noise is relatively small. These findings are consistent with the design principles of our benchmark.
Table 2: Benchmarking on ViLP. Please refer to the left text for symbol definitions. †  indicates the model often fails to follow the instructions.
Model ViLPF ViLPP
Score Prior Score Prior
Baseline
Human 98.33 99.67 98.67 96.67
GPT-4o (text only) 0.0 92.33 0.17 71.33
API call only
GPT-4o 66.17 91.00 56.00 87.67
GPT-4V 57.67 88.33 38.33 85.33
GPT-4o-Mini 57.67 89.00 46.67 84.67
Claude-3.5-Sonnet 70.00 84.33 59.33 86.67
Claude-3-Opus 59.17 74.00 43.00 82.67
Claude-3-Sonnet 48.83 83.67 40.33 81.33
Claude-3-Haiku 43.67 82.67 34.83 82.33
Gemini-1.5-Pro 60.50 79.33 48.00 83.00
Gemini-1.5-Flash 54.50 83.33 69.17 79.67
Open weights
Llama-3.2-Vision-11B 67.33 76.67 61.17 79.33
Llama-3.2-Vision-90B 64.00 91.67 63.17 83.33
MolmoE-1B 48.67 57.33 47.83 69.00
Molmo-7B-O 57.83 60.67 47.33 76.33
Molmo-7B-D 54.5 69.00 46.17 72.33
Molmo-72B 60.33 85.00 47.17 82.33
Qwen2-VL-7B 50.50 83.00 48.67 80.33
Qwen2-VL-72B 56.50 92.33 53.83 83.00
InternVL2-8B 47.00 66.67 43.00 75.00
InternVL2-76B 42.67 47.67 50.84 74.33
LLaVA-1.5-7B 29.67 71.33 37.67 65.67
LLaVA-1.5-13B 35.33 81.00 41.50 73.67
Cambrian-1-8B 8.67 43.67 32.50 63.67
LLaVA-OneVision-7B 54.17 82.33 49.67 75.00
LLaVA-OneVision-72B 1.67 3.00 5.22 11.67
Figure 4: Before and after removing distactor facts.
Are our QIAs aligned with the learned priors of VLMs? We evaluated GPT-4o’s performance on our questions without providing any images, i.e., GPT-4o (text only) in Table 2. Additionally, we eliminate image-related phrases from the prompt during this text-only inference, such as removing expressions like “as shown in the image". Remarkably, even without any visual content, the model was able to correctly answer 92.33% of QIAprior when distractor facts were included, demonstrating that our questions are highly suggestive for VLMs, allowing them to answer even without visual input. Once removing the distractor facts, the performance decrease to 71.33%, as these facts implicitly pointed to the correct answers in the fact statements. It is noteworthy that some powerful models like Claude-3-Opus and several open-source VLMs fail to outperform GPT-4o (text-only) performance on QIAprior, even when provided with images. For QIAtest, the success rate of GPT-4o (text-only) dropped dramatically to near zero (0% & 0.17%), indicating that QIAtest are far outside the distribution typically encountered by VLMs or LLMs. These observations further validate the design principles of our benchmark: QIAprior can be easily inferred using text alone, while QIAtest is specifically designed to challenge language-based reasoning, requiring reliance on visual cues to arrive at correct answers.
How do VLMs perform on our benchmark? Although our benchmark questions are distinguishable for humans, they are challenging for VLMs. Even the most advanced VLM models like GPT-4o, have a clear performance gap (more than 25%) compared to humans’ performance on QIAtest (Score in ViLPF), indicating the difficulty of these questions for VLMs. GPT-4o achieves above 91% accuracy on QIAprior (Prior), which is quite close to human performance. While its Score on QIAtest is still low (66.17% versus 98.33%), demonstrating the difficulty of these questions for VLMs. We also provide some qualitative examples in Figure 3, showcasing results from both leading commercial and open-source models, including GPT-4o, Claude-3.5-Sonnet, Gemini-1.5-Pro, and Llama-3.2-Vision-90B. They encounter challenges when tackling these cases of our ViLP. Notably, it is encouraging to see that some open-source models achieved over 60% accuracy on ViLPF-QIAtest, with performance nearing that of their commercial counterparts, including Llama-3.2-Vison and Molmo-72B.
Do distractor facts really distract? In ViLPF setting, we add a distractor fact before the question and explicitly instruct the model to answer with one word. Since these facts implicitly suggest incorrect answers for QIAtest, we expected this change to make the questions more suggestive and lower the ViLPFScore, as the distractors would mislead the VLMs.
But is this true in practice? Interestingly, we find that strong models like GPT-4o are not misled by these distractor facts, despite their performance is still significantly lagged behind human accuracy. Not only do these models avoid being misled, but surprisingly, they also leverage the distractor factsto improve their answers, leading to a substantial and consistent performance boost when the facts are included. We illustrate some examples of this behavior in Figure 4. When the distractor fact is provided, GPT-4o performs better on QIAtest compared to when it is not. This result is counterintuitive, as the model performs better when its input prompt includes misleading information. We hypothesize that the distractor fact may help the model infer the focus of the question, thereby narrowing the search space and ultimately leading to higher accuracy.
For models with reasonable instruction-following abilities but limited visual reasoning, like LLaVA-1.5-13B (Liu et al., 2023b), the distractor facts behave as expected, often leading the model to incorrect conclusions based solely on textual information. This leads to a significant performance drop on ViLPF-Score when distractor facts are present, simultaneously boosting performance on ViLPF-Prior, the language-prior answer. For instance, as shown in Figure 4, with including distractor facts, LLaVA-1.5-13B consistently predicts the distractor fact as the answer. However, once the distractors are removed, it can then make the correct prediction.
In contrast, for models with weaker instruction-following abilities, like Cambrian-8B (Tong et al., 2024), the inclusion of distractor facts significantly increases the difficulty of adhering to explicit instructions, such as providing a single-word answer. Specifically, when facts are provided, Cambrian-8B fails to follow the instruction in 62% of the questions, compared to 30% without facts (nearly a 2x increase). Upon manual review of these failures, we find that approximately 59% of the responses that deviate from the instruction are still contextually correct, leading to an adjusted accuracy of 47.92%. A similar issue is observed with LLaVA-OneVision-72B (Li et al., 2024b), which consistently generates sentences with analysis, when explicitly prompted with “please answer this question with one word". This pattern highlights a concerning trend: focusing on improving performance on well-established benchmarks may come at the cost of basic instruction-following abilities, ultimately limiting the practical utility of these models in real-world applications.
Figure 5: Comparison of benchmark scores under different image transformations. Solid line and dotted line refer to ViLPF-Score and ViLPF-Prior, respectively.
How image transformations affect the results? We also investigate how image transformations, including resizing, Gaussian blur, and pixelation, affect ViLPperformance. The results, shown in Figure 5, reveal that the ViLPF-Score rapidly decreases as the severity of the transformations (x-axis) increases, while the ViLPF-Prior score remains around 50%. Interestingly, GPT-4o, when using degraded images, performs worse in ViLPF-Prior than when no images are used, i.e., GPT-4o (text only) in Table 2.
5.1 Image DPO
In this section, we examine our proposed Image-DPO (Section 4.1) on both the ViLP  benchmark and general VQA benchmarks. Specifically, for self-generating question-image-answer (QIA) pairs, we use COCO (Lin et al., 2014), Text2VQA (Singh et al., 2019b), and Visual Genome (Krishna et al., 2017) as our seed datasets. Based on these seed datasets, we instruct the VLM to leverage pretrained image generation and editing models (Liu et al., 2023d; Podell et al., 2023; Brooks et al., 2023b) to generate new images and corresponding QA pairs. After generating the QA pairs, we corrupt the images either semantically or at the pixel level, while keeping the QA pairs unchanged, to create the good and bad QIA pairs used for Image DPO training. Details of our image generation method are provided in the Appendix C and training details are in the Appendix C.2
Table 3: Comparisons of Image-DPO on ViLP.
Model ViLP F ViLP P
Score Prior Score Prior
LLaVA-1.5-7B 29.67 71.33 37.67 65.67
LLaVA-1.5-7B + CSR 27.50 64.33 40.84 63.67
LLaVA-1.5-7B + RLHF-V 29.50 75.00 36.33 65.33
LLaVA-1.5-7B + Text-DPO 31.34 71.67 37.83 65.67
LLaVA-1.5-7B + Image-DPO 34.17 77.67 38.67 67.33
For comparison, we introduce another baseline, Text-DPO, which follows the same question-image-answer generation process as our Image-DPO but applies LLM self-rewarding techniques as outlined in (Yuan et al., 2024). In Text-DPO, good and bad pairs are created by sampling positive and negative answers generated by the VLM, while keeping the image and question unchanged. The difference between Image-DPO and Text-DPO is illustrated in Figure 6. The green box corresponds to Image-DPO, where multiple corrupted images are generated through semantic alterations, Gaussian blurring, and pixelation, while the question and answer remain constant. In contrast, Text-DPO (purple-box) maintains the same image and generates multiple answers, each associated with a rating. It is worth noting that while Text-DPO shares similarities with other VLM self-rewarding techniques (Zhou et al., 2024a; Deng et al., 2024; Zhou et al., 2024c; Wang et al., 2024b; a), a key distinction is our use of pre-trained image generation models to create diverse images.
Table 4: Effectiveness of Image-DPO on General VQA benchmarks. The proposed Image-DPO could consistently improve the performance of various VLMs across general VQA benchmarks.
[…]
Cambrian-8B 8.67 32.50 44.6 47.9 19.4 14.5 4.7 51.4 11.3
Cambrian-8B + Image-DPO 11.50 $\uparrow$ 2.83 35.17 $\uparrow$ 2.67 45.68 $\uparrow$ 1.08 49.45 $\uparrow$ 1.55 20.11 $\uparrow$ 0.71 14.3 $\uparrow$ -0.2 4.8 $\downarrow$ 0.1 52.7 $\uparrow$ 1.3 12.0 $\uparrow$ 0.7
LLaVA-1.5-7B 29.67 37.67 37.7 43.8 12.7 49.1 14.8 31.1 58.6
LLaVA-1.5-7B + Image-DPO 34.17 $\uparrow$ 4.5 38.67 $\uparrow$ 1 40.68 $\uparrow$ 2.98 46.29 $\uparrow$ 2.49 14.95 $\uparrow$ 2.25 45.4 $\uparrow$ -3.7 12.4 $\uparrow$ -2.4 33.8 $\uparrow$ 2.7 59.65 $\uparrow$ 1.05
LLaVA-1.5-13B 35.33 41.50 39.6 44.6 14.8 48.30 14.10 36.10 61.9
LLaVA-1.5-13B + Image-DPO 38.17 $\uparrow$ 2.84 42.5 $\uparrow$ 1 42.68 $\uparrow$ 3.08 47.37 $\uparrow$ 2.77 17.16 $\uparrow$ 2.36 42.6 $\uparrow$ -5.7 11.6 $\uparrow$ -2.5 37.6 $\uparrow$ 1.5 62.32 $\uparrow$ 0.42
Figure 6: Image-DPO v.s. Text-DPO.
For comparison, we also report the results of CSR (Zhou et al., 2024b), a VLM self-improvement method that similarly constructs good and bad answers to form training data for DPO. Specifically, we use the publicly available checkpoint of the CSR model. Additionally, we train a model using RLHF-V (Yu et al., 2023a), a supervised fine-tuning method based on human-annotated data. All models are trained on the LLaVA-7B architecture due to computational constraints. The results, presented in Table 3, show that our Image-DPO method achieved the highest performance in ViLPF-Score, outperforming Text-DPO, CSR (Zhou et al., 2024b), and the RLHF-V model based on human-annotated data (Yu et al., 2023a). On ViLPP-Score, the Image-DPO method achieved the second-best performance.
Besides, we evaluate the proposed Image-DPO algorithm across three VLM models—Cambrian-8B, LLaVA-1.5-7B, and LLaVA-1.5-13B—using several popular VLM benchmarks that focus on different aspects, including compositionality(NaturalBench (Li et al., 2024a)) hallucinations (CHAIR (Rohrbach et al., 2018)), visual conversation (MM-Vet (Yu et al., 2023c)), and visual reasoning, SEED-Bench (Li et al., 2023a)). The results, presented in Table 4, show consistent performance improvements across both datasets and models, further demonstrating the effectiveness of our Image DPO method.
6 Conclusion
In conclusion, we present the ViLPbenchmark to probing the challenge of language bias in Vision-Language Models. By utilizing advanced image generation models and designing questions that demand visual cues for accurate responses, our benchmark consists of images that challenge language priors, revealing the limitations of current VLMs, which tend to rely heavily on text. Our method, which incorporates self-generated VQA pairs and image corruption for training, has shown promising improvements in enhancing visual reliance, as evidenced by performance gains on models like LLaVA-v1.5 and Cambrian.
7 Acknowledgment
This work was supported by the LG AI Research grant. We also extend our gratitude to the OpenAI Researcher Access Program for providing the credits used to access OpenAI’s APIs. Finally, we greatly appreciate Chris Rockwell for his contributions to our human evaluation studies.
Appendix
[…]
Appendix A More data samples of ViLP
Figure 7: Randomly sampled data from ViLP.
Figure 8: Randomly sampled data from ViLP.
Appendix B Image-DPO Mathematical Details
Figure 9: Gradients difference between Image-DPO and Text-DPO. For Text-DPO (Rafailov et al., 2024), the model receives positive gradients (green arrows) for the preferred answer $A_{w}$ and negative gradients $A_{l}$ (red arrows) for the dispreferred answer. In contrast, our proposed Image-DPO approach applies positive gradients when the preferred image $I_{w}$ is input and negative gradients for the dispreferred image $I_{l}$, both based on the same output answer.
In this section, we give the complete proof of Image DPO. Simlarly to DPO, we start from the objectiveness of RLHF and then derivate its variant for image dpo.
B.1 RLHF for VLM
SFT Given question $Q$, answer $A$ and image $I$, we can train a SFT model $\pi^{SFT}$ with supervised learning on high-quality data. As the SFT model is a language generation model, it is still a model modeling the text outputs with question and image. $\pi^{SFT}(A|Q,I)$
Reward Modeling Phase In this stage, we construct a static dataset of comparisons $\mathcal{S}=\{A^{i},Q^{i},I_{w}^{i},I_{l}^{i}\}$, and we present the QIA pairs $(Q,I_{w},A)$, $(Q,I_{l},A)$ to human for preference.
Following the idea of RLHF, the preference are assumed to be obtained from a a latent reward function $r^{*}(Q,I,A)$ which are not tractable, and we we use BT model to represent the preference distribution $p^{*}$ as:
$p^{*}((Q,I_{w},A)\succ(Q,I_{l},A))=\frac{\exp\left(r^{*}\left(Q,I_{w},A\right)\right)}{\exp(r^{*}(Q,I_{w},A)+\exp(r^{*}(Q,I_{l},A))}$ (5)
Now given the human labeled preference, we can try to optimize a reward model $r_{\phi}$ to estimate $r^{*}$ by using maximum likelihood. Framing this as a binary classification, we can have this negative log-likelihood loss:
$\mathcal{L}_{R}(r_{\phi},\mathcal{S})=-\mathbb{E}_{(A,Q,I_{w},I_{l})\sim\mathcal{S}}\left[\log\sigma(r_{\phi}(Q,I_{w},A)-r_{\phi}(Q,I_{l},A))\right]$ (6)
Here $\sigma$ is a logistic function. Basically, this reward function gives score jointly considering image, question and image quality.
RL Fine-Tunning Phrase
During the RL phase, the learned reward function is used to provide feeback to the VLM model. Following DPO paper, the optimization is formulated as :
$\max_{\pi_{\theta}}\mathbb{E}_{(Q,I)\sim\mathcal{S},A\sim\pi_{\theta}(A|Q,I)}\left[r_{\phi}(Q,I,A)\right]-\beta\mathbb{D}_{\text{KL}}\left[\pi_{\theta}(A|Q,I)\|\pi_{\text{ref}}(A|Q,I)\right],$ (7)
where $\beta$ is a parameter controlling the deviation from the base reference policy $\pi_{\text{ref}}$, namely the initial SFT model $\pi^{\text{SFT}}$. Due to the discrete nature of language generation, this object is also not differentiable and is typically optimized with reinforcement learning.
B.2 Image DPO and RLHF
According to the DPO paper, a straightforward optimal solution to the KL-constrainted reward function maximization object in Eq. 6 is:
$\pi_{r}(A|Q,I)=\frac{1}{Z(Q,I)}\pi_{\text{ref}}(A|Q,I)\exp(\frac{1}{\beta}r(Q,I,A))$ (8)
where $Z(Q,I)=\sum_{A}\pi_{\text{ref}}(A|Q,I)\exp(\frac{1}{\beta}r(Q,I,A)$ is a partition function. Here $r$ should be any reward function, which makes $Z$ hard to tract. We provide the proof of this step in B.3.
Taking the logarithm of both side, and with some algebra, we get
$r(Q,I,A)=\beta\frac{\pi_{r}(A|Q,I)}{\pi_{\text{ref}}(A|Q,I)}+\beta\log Z(Q,I)$ (9)
This parametrization could be applied to ground-truth reward $r^{*}$ and the corresponding optimal model $\pi^{*}$.
The BT model with the optimal policy is
$p^{*}((Q,I_{1},A)\succ(Q,I_{2},A))=\frac{\exp\left(r^{*}(Q,I_{w},A)\right)}{\exp\left(r^{*}(Q,I_{w},A)\right)+\exp\left(r^{*}(Q,I_{l},A)\right)}$ (10)
We plug Eq. 9 into the BT model, we have:
$p^{*}((Q,I_{1},A)\succ(Q,I_{2},A))$ $=\frac{\exp\left(\beta\log\frac{\pi^{*}(A|Q,I_{w})}{\pi_{\text{ref}}(A|Q,I_{w})}+\beta\log Z(Q,I_{w})\right)}{\left(\beta\log\frac{\pi^{*}(A|Q,I_{w})}{\pi_{\text{ref}}(A|Q,I_{w})}+\beta\log Z(Q,I_{w})\right)+\left(\beta\log\frac{\pi^{*}(A|Q,I_{l})}{\pi_{\text{ref}}(A|Q,I_{l})}+\beta\log Z(Q,I_{l})\right)}$
$=\frac{1}{1+\exp(\beta\log\frac{\pi^{*}(A|I_{l},Q)}{\pi_{\text{ref}}(A|I_{l},Q)}-\beta\log\frac{\pi^{*}(A|I_{w},Q)}{\pi_{\text{ref}}(A|I_{w},Q)}+\beta\log Z(I_{l},Q)-\beta\log Z(I_{w},Q))}$
$=\sigma\left(\exp(\beta\log\frac{\pi^{*}(A|I_{l},Q)}{\pi_{\text{ref}}(A|I_{l},Q)}-\beta\log\frac{\pi^{*}(A|I_{w},Q)}{\pi_{\text{ref}}(A|I_{w},Q)}+\beta\log Z(I_{l},Q)-\beta\log Z(I_{w},Q))\right)$
Now we have the probability of human preference data in terms of the optimal policy rather than the reward model, we can formulate a maximum likelihood objective for a policy $\pi_{\theta}$. Our policy objective is :
$\mathcal{L}(\pi_{\theta};\pi_{\text{ref}})=\mathbb{E}_{(Q,A,I_{w},I_{l})\sim\mathcal{S}}\left[-\log\sigma\left(\beta\log\frac{\pi_{\theta}(A|I_{w},Q)}{\pi_{\text{ref}}(A|I_{w},Q)}-\beta\log\frac{\pi_{\theta}(A|I_{l},Q)}{\pi_{\text{ref}}(A|I_{l},Q)}+\beta\log Z(I_{w},Q)-\beta\log Z(I_{l},Q)\right)\right]$ (11)
As $f(x)=-\log\sigma(x)$ is a convex function ($\sigma$ is the sigmoid function), we can apply Jensen’s inequality $f(\frac{1}{2}x+\frac{1}{2}y)\leq\frac{1}{2}f(x)+\frac{1}{2}f(y)$:
$\mathcal{L}(\pi_{\theta};\pi_{\text{ref}})\leq\mathbb{E}\left[-\frac{1}{2}\log\sigma\left(2\beta\log\frac{\pi_{\theta}(A|I_{w},Q)}{\pi_{\text{ref}}(A|I_{w},Q)}-2\beta\log\frac{\pi_{\theta}(A|I_{l},Q)}{\pi_{\text{ref}}(A|I_{l},Q)}\right)-\frac{1}{2}\log\sigma\left(2\beta\log Z(I_{w},Q)-2\beta\log Z(I_{l},Q)\right)\right]$ (12)
As $\log\sigma\left(Z(I,Q)\right)$ is not a function of $\pi_{\theta}$, the above objective is equivalent to the below Eq.13, where $\alpha=2\beta$. It is the same as our objective listed in Eq.4 of the main paper.
$\mathcal{L}(\pi_{\theta};\pi_{\text{ref}})\leq-\mathbb{E}_{(Q,I_{w},I_{l},A)\sim S}\left[\log\sigma(\alpha\frac{\pi_{\theta}(A\mid Q,I_{w})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{w})}-\alpha\frac{\pi_{\theta}(A\mid Q,I_{l})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{l})})\right]$ (13)
In this sense, our optimization objective Eq.4 in main paper are optimizing the upper bound of RLHF, i.e., Eq.7.
B.3 Deriving the Optimum of the KL-Constrained Reward Maximization Objective
In this appendix, we will derive Eq.8. Similarly to Eq.7, we optimize the following objective:
$\max_{\pi}\mathbb{E}_{(Q,I)\sim\mathcal{S},A\sim\pi}\left[r(Q,I,A)\right]-\beta D_{\mathrm{KL}}\left[\pi(A|Q,I)\|\pi_{\mathrm{ref}}(A|Q,I)\right]$ (14)
under any reward function $r(Q,I,A)$, reference model $\pi_{\mathrm{ref}}$, and a general non-parametric policy class. We now have:
$\max_{\pi}\mathbb{E}_{(Q,I)\sim\mathcal{S},A\sim\pi}\left[r(Q,I,A)\right]-\beta D_{\mathrm{KL}}\left[\pi(A|Q,I)\|\pi_{\mathrm{ref}}(A|Q,I)\right]$ (15)
$=\max_{\pi}\mathbb{E}_{(Q,I)\sim\mathcal{S}}\mathbb{E}_{A\sim\pi(A|Q,I)}\left[r(Q,I,A)-\beta\log\frac{\pi(A|Q,I)}{\pi_{\mathrm{ref}}(A|Q,I)}\right]$
$=\min_{\pi}\mathbb{E}_{(Q,I)\sim\mathcal{S}}\mathbb{E}_{A\sim\pi(A|Q,I)}\left[\log\frac{\pi(A|Q,I)}{\pi_{\mathrm{ref}}(A|Q,I)}-\frac{1}{\beta}r(Q,I,A)\right]$
$=\min_{\pi}\mathbb{E}_{(Q,I)\sim\mathcal{S}}\mathbb{E}_{A\sim\pi(A|Q,I)}\left[\log\frac{\pi(A|Q,I)}{\frac{1}{Z(Q,I)}\pi_{\mathrm{ref}}(A|Q,I)\exp\left(\frac{1}{\beta}r(Q,I,A)\right)}-\log Z(Q,I)\right]$
where we have the partition function:
$Z(Q,I)=\sum_{A}\pi_{\mathrm{ref}}(A|Q,I)\exp\left(\frac{1}{\beta}r(Q,I,A)\right)$ (16)
Observe that the partition function depends solely on $(Q,I)$ and the reference policy $\pi_{\mathrm{ref}}$, and is independent of the policy $\pi$. We can now define the Equation 8.
$\pi^{*}(A|Q,I)=\frac{1}{Z(Q,I)}\pi_{\mathrm{ref}}(A|Q,I)\exp\left(\frac{1}{\beta}r(Q,I,A)\right),$ (17)
Appendix C Details in Image-DPO data generation and training
Our image-DPO data generation pipeline consists of two stages. In the first stage, we leverage the VLM we aim to enhance to perform self-guided data generation with the aid of pre-trained image generative models. This stage produces a large number of new question-image-answer (QIA) triplets. In the second stage, we apply three types of image corruptions—Gaussian blurring, pixelization, and semantic editing—to generate good-bad QIA pairs, denoted as $I_{w}$ (good) and $I_{l}$ (bad).
Details of the hyperparameters used in the experiments are provided at the end of this section.
C.1 VLM self-guided data generation
Figure 10: Overview of our data generation pipeline. We begin with an image only, from which instructions are derived using a VLM. These instructions guide the creation of a new image or the modification of the existing one. The generated image is then processed by the VLM to generate QA pairs. Both the QA pair and the image are subsequently input back into the VLM to assess the quality of the answers. No human-written in-context examples are used throughout this process.
As illustrated in Figure 10, our data generation process begins by utilizing VLMs to suggest modifications or draw inspiration for input images without relying on any in-context examples. The used text prompt is shown in Figure 12. Subsequently, pre-trained models such as Stable Diffusion XL Podell et al. (2023), Instruct-Pix2Pix Brooks et al. (2023b), and Grounded-SAM Ren et al. (2024) are employed to either modify existing images or generate entirely new ones.
The altered or newly created images, along with the instructions that guided their generation, are then used by the same VLMs to produce corresponding question-answer pairs (QAs) based on the text prompt shown in Figure 13. An example of this process is provided in Figure 11. Importantly, all instructions, tool selections, and QA generation are autonomously handled by the same VLM we aimed to improve.
In particular, Grounded-SAM requires the VLM to specify the object to be modified before generating images. To facilitate this, we use an additional text prompt (Figure 14) after the VLM generates the initial instructions (the pink region of Figure 11).
To provide a better understanding of our generated QIAs, we randomly sampled and listed some examples of the generated QIA data, as shown in Figures 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, and 29.
Figure 11: Randomly sampled generation examples. Our proposed data generation pipeline utilizes pretrained model to generate diverse new images from existing seed image datasets.
Given this image, please suggest a range of creative edits, tasks, or transformations that could be applied using advanced image processing tools. These tasks may include artistic transformations, vivid color adjustments, object detection and modification, or completely creating a new image inspired by the original. Specify which tool would be best suited for each task, choosing from Stable Diffusion for image generation, InstructPix2Pix for image modification, or GroundingDINO for object modification. Your recommendations should help in understanding the potential of the image and exploring creative possibilities. Expected Response Format: Item Number: 1 Tool Used: [Specify the tool - Stable Diffusion or InstructPix2Pix or GroundingDINO] Text Prompt for Processing: [Detailed description of the task or transformation to be performed. For image generation, please provide complete description based on the understanding of the provided images, since we only feed text prompt for this task.] Item Number: 2 Tool Used: [Specify the tool - Stable Diffusion or InstructPix2Pix or GroundingDINO] Text Prompt for Processing: [Detailed description of the task or transformation to be performed. For image generation, please provide complete description based on the understanding of the provided images, since we only feed text prompt for this task.] Item Number: 3 Tool Used: [Specify the tool - Stable Diffusion or InstructPix2Pix or GroundingDINO] Text Prompt for Processing: [Detailed description of the task or transformation to be performed. For image generation, please provide complete description based on the understanding of the provided images, since we only feed text prompt for this task.]
Figure 12: The prompt for instruction generation. We ask the VLM to generate instructions for using pre-trained image models.
Given this image, could you please generate a series of insightful and diverse question-answer pairs based on the image and its descriptions? We are interested in exploring various facets of the image, including: - Holistic styles and layouts: Questions that analyze the overall design, style, and layout of the image. - Object-specific details: Questions that delve into particular elements or objects within the image, discussing their characteristics or functions. - Background context: Questions that speculate about the background story or the setting of the image. - Overall themes: Questions that interpret the thematic elements and messages portrayed in the image. We encourage creative and thought-provoking questions that extend beyond the basics. Please generate questions that cover a broader range of observations and insights drawn from the image. Each question should be followed by a comprehensive answer, providing depth and context. Expected Multiple Response Format: Item Number: 1 Question: [Propose a unique and insightful question based on the descriptions and the images.] Answer: [Provide a comprehensive answer to the proposed question.] Item Number: 2 Question: [Propose a unique and insightful question based on the descriptions and the images.] Answer: [Provide a comprehensive answer to the proposed question.] Please ensure each question-answer pair is well-defined and informative. Please provide at least 5 question-answer pairs based on the input provided.
Figure 13: The prompt for single-image QAs. We ask the VLM itself to generate single-image QAs based on the generated images by pre-trained models.
Analyze the provided image and its accompanying modification instruction to identify the removed object description, the new object description, and the new image description. Modification Instructions: <Text Prompt for Processing> Expected Multiple Response Format: Item Number: 1 Removed Object Description: [Brief description of the object to be detected and removed] New Object Description: [Description of a new, different object to replace the removed one] New Image Description: [Description of the image after each object’s removal, focusing on changes and remaining elements] Item Number: 2 Removed Object Description: [Brief description of the object to be detected and removed] New Object Description: [Description of a new, different object to replace the removed one] New Image Description: [Description of the image after each object’s removal, focusing on changes and remaining elements]
Figure 14: The prompt for instruction generation of Grounded-SAM. We ask the VLM to generate designated instructions to use Grounded-SAM.
C.2 Image DPO data preparation and training details
This section details the construction of good-bad question-image-answer (QIA) pairs ($I_{w}$, $I_{l}$) based on the QIAs generated by the pipeline described in Appendix C.1. In brief, the data generation pipeline outlined in Appendix C.1 utilizes VLMs in conjunction with pre-trained image models to generate or modify images and create corresponding question-answer pairs. This process results in a collection of QIA triplets, as illustrated in the Figures 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, and 29.
After generating the QIA triplets, we apply three image corruption methods—Gaussian blurring, pixelization, and semantic editing—to create good-bad QIA pairs for ImageDPO training (Section 4.1), while keeping the QA components unchanged.
For Gaussian blur, we use a kernel size of 40 for Cambrian and 80 for LLaVA, as the larger kernel size showed better performance for LLaVA. For pixelization, we apply block sizes of 32 for Cambrian and 64 for LLaVA. For these two types of corruption, we utilize data generated by all three pre-trained models: Stable Diffusion, Instruct-Pix2Pix, and Grounded-SAM.
However, for semantic editing, we rely solely on data from Grounded-SAM, which modifies a single object in the image while leaving the rest unchanged. The object to be altered and the type of modification are determined by the VLMs based on the text prompt (Figure 14). For instance, the VLM decides to add a headset to the chef, as shown in Figure 15. Grounded-SAM then modifies the image accordingly by adding the headset, after which the VLM generates a single QA pair based on the text prompt (Figure 13). Note that, we retain only images where the altered region covers more than 10% of the image size, as this ensures a significant and distinct change.
For model training, we constructed datasets comprising 800k and 400k DPO pairs to fine-tune LLaVA (7B and 13B) and Cambrian-8B, respectively. Batch sizes are set to 112 for LLaVA-7B, 80 for LLaVA-13B, and 8 (with 4 gradient accumulation steps) for Cambrian-8B. We employ LoRA with a rank of 128, an alpha of 256, and a learning rate of 5e-7, training each model for 2 epochs. The GPUs we used are 8-L40S.
Figure 15: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 16: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 17: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 18: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 19: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 20: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 21: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 22: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 23: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 24: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 25: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 26: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 27: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 28: Randomly sampled single-image QA generated by our data generation pipeline.
Figure 29: Randomly sampled single-image QA generated by our data generation pipeline.
Appendix D More details and comparisons of our benchmarks
D.1 Our benchmark data generation
Our proposed dataset introduces QIA triplets meticulously designed to challenge state-of-the-art Vision-Language Models (VLMs) by mitigating language priors. The construction process combines human-guided and automated efforts to ensure quality and alignment across all components.
Question-Answer (QA) Generation: Questions and answers are partially authored by humans, adhering to the design principles outlined in Section 3.1. Additionally, candidate QA pairs are generated using models like OpenAI-O1 and Claude-3.5-Sonnet with carefully crafted prompts. One example prompts shown in Figure 30. These candidates undergo rigorous human review, where they are refined or filtered to meet our quality standards.
Image Generation: To align with the QA pairs, descriptive and diverse image prompts are first generated using GPT-4 (see Figure 31). These prompts are then input into advanced image generation models, such as FLUX and DALL-E 3, to produce candidate images. Multiple images are generated per QA pair, and human reviewers carefully select the most suitable image or request re-generation to ensure alignment with the QA context.
Human Review and Testing: At every stage, human reviewers rigorously evaluate the generated outputs to ensure quality, clarity, and challenge level. In addition to filtering out low-quality or insufficiently challenging triplets, we also test the QIAs with users to confirm that they remain intuitive for humans while being difficult for VLMs.
Overall, the complexity of creating our dataset results in a high average cost of generating a single QIA triplet in ViLP, amounting to approximately \$7.50, excluding human labor costs.
In the below, I try to propose questions along with three answers where the first answer is corresponding to the question text directly, while the other two are usual and counter-intuitive, which could lead to wrongs of VLMs. Please help me generate more Question-3 answer pairs, which are different from what I have provided. - All the potential answers should a single world. - Help me generate a format where I can direct copy paste into Goole Sheet. Also, please a ; between question and each answers. - Please be very creative and different from my provided examples - the answer 2 & 3 should be very diverse and different compared to answer 1. - Every question contains a statement at the beginning which consists of the answer1 as part of it. - Please understand the principles and generate the QA very different from my provided examples Some Examples: - A screwdriver is used for tightening screws. From the image, which tool is used to turn screws? Screwdriver Hammer Scissors - A pen is a tool used for writing. Which object in the image is used to write on paper? Pen Hammer Shoe - Clocks are used to measure time. Can you identify the item in the image that is used to measure time? Clock Spoon Candle - A violin has four strings and is played using a bow. According to the image, which musical instrument is being played with a bow? Violin Guitar Saxophone - Camels have humps. Which animal in the image stores fat in its humps? Camel Horse Tiger - Honey is made by bees. Which insect in the image produces honey? Bee Ant Dragonfly - An anvil is a tool used by blacksmiths. What object in the image is used by blacksmiths to forge metal? Anvil Fork Wrench - A gavel is used by judges in court. From the image, which object symbolizes judicial authority? Gavel Hammer Wrench - A syringe is used to inject medicine. From the image, which tool is used for administering injections? Syringe Scissor Drill - An anchor keeps a ship steady in the water. From the image, which item prevents boats from drifting? Anchor Spoon Toothbrush - A chainsaw is a power tool for cutting wood. What device shown is typically used by lumberjacks to fell trees? Chainsaw Blender Stapler
Figure 30: One prompt we used for potential QAs designs of ViLP
Task: Using the provided question and possible image-based answers, generate detailed text prompts for image generation. Each image prompt should reflect the question’s context and incorporate one of the image-based answers. Question: Question Image-based Answer:[Answer1, Answer2, Answer3] For each possible image-based answer, create an image prompt that describes what the image might look like based on the question. Please be creativity. For example, if the question asks who is using this mop to clean the floor in the picture? and the answer is eraser. The image prompt should really describe the image of an eraser uses a mop to clean the floor. Format the output strictly as a JSON list, like this example: [ "prompt1": "Image Generation Prompt text here", "prompt2": "Image Generation Prompt text here", "prompt3": "Image Generation Prompt text here", ]
Figure 31: The prompt we used for generating text-prompt for image generation.
D.2 Comparisons to other datasets
Here we compare our benchmark to other popular benchmarks. All these compared datasets are well-designed and impactful. However, they have different evaluation focus compared to ours, from high-level design principles to low-level formats.
D.2.1 Compare to Winoground Thrush et al. (2022)
Winoground focuses on vision-linguistic compositional reasoning by presenting models with two images and two captions containing identical sets of words arranged differently. The goal is to match images to captions based on compositional structures in the text and visual content (as detailed in the Introduction and Sec 3.1 of the referenced paper). However, the captions in Winoground do not challenge language priors or introduce out-of-distribution information. Both captions remain consistent with common linguistic expectations, and no explicit misleading information is provided to test resistance to language biases.
Figure 32: Winoground Data Example. Our benchmark is different from Winoground, as Winoground focuses on vision-linguistic compositional reasoning. Both captions and images are normal and satisfy common linguistic expectations and common sense.
Qualitative Comparison As shown in Figure 32 consisting of Winoground examples, both captions and images are normal and satisfy common linguistic expectations and common sense. The evaluation focus is on whether or not the model can tell the compositional difference between the two images and two captions, and match them correctly.
Quantitative Comparison Both ViLP  and Winoground benchmarks include paired textual information in their settings. In our benchmark, we have ViLP  Prior QAs and ViLP  Score QAs, which share the same question but differ in their answers. In Winoground, for each example, there are two captions, and the task is to match the correct image to the correct caption.
To demonstrate the differences, we use GPT to evaluate the commonness of these paired textual components. Specifically, GPT-4o rates the oddity of scenarios described in texts on a scale from 1 (very rare) to 10 (very common), and these scores are compared.
In our benchmark, Prior QAs scored 9.37, indicating answers designed to align with language priors are easily guessed and highly common. Score QAs scored 1.65, showing that these QA pairs are significantly rare, which are difficult to infer without the corresponding visual information. Notably, Prior and Score QAs share the same question but differ in their answers. The stark difference in scores demonstrates our injection of strong language priors, used to test the model’s susceptibility to linguistic distractions. For comparison, Winoground’s two captions yielded scores of 8.05 and 8.08, revealing two major differences:
Both captions align perfectly with language priors (at least according to GPT-4o), indicating that Winoground does not challenge language priors or evaluate out-of-distribution scenarios The minimal score difference between the two captions indicates no significant variance in language priors, as exploring how VLM models behave under varying language priors is not within the scope of this benchmark. In contrast, this is precisely the focus of our benchmark.
D.2.2 Compare to Whoops! Bitton-Guetta et al. (2023a)
Whoops! is designed to evaluate a model’s ability to detect weirdness in images, focusing on tasks where images depict unusual or nonsensical scenarios. It highly emphasizes common sense reasoning: models are expected to recognize visual elements and then reason the subtle contrast between these elements to identify the oddity. For example, for A lit candle is sitting inside a tightly sealed glass jar example in the homepage, models are expected to realize that A candle needs a constant supply of oxygen to burn, which does not exist in a sealed bottle, so it is unlikely to see a burning candle inside a sealed bottle. This benchmark emphasizes common-sense reasoning instead of defying language priors.
Figure 33: Example of Whoops! Dataset.Whoops! also has creative images. While unlike ours, its questions are common questions without strong language priors.
Qualitative Comparison
Although its images are creative and out of distribution like ours, this benchmark does not focus on using language priors to test the model’s susceptibility to linguistic distractions as our benchmark does. Especially, in their QA mode(which is the same task as our benchmark), the questions are normal questions without strong language priors like ours. Some examples can be found in Figure 33. Also, it uses a format of open-ended questions. This format allows higher freedom in QA, while adding ambiguity and even offs in the answers.
Quantitative Comparison
Unlike Winoground and our benchmark, Whoops! does not provide control groups or textual components for comparison. To measure language priors, we analyze the suggestiveness of questions by evaluating how determined and confident GPT-4o answers them (without visual context). A more suggestive question is expected to yield more determined and confident answers, while a less suggestive question is expected to yield more diverse answers. For this, we calculate the number of unique answers given by GPT-4o over five attempts while setting its temperature to 1.0 for higher randomness. Semantic differences were normalized to exclude synonyms.
Interestingly, we find questions from Whoops! yield 2.58 unique answers (in 5 attempts) on average with std 1.48. For our benchmark, without facts, GPT-4o gives 1.53 unique answers (in 5 attempts) on average with std 0.94. With facts, GPT-4o gives 1.10 unique answers (in 5 attempts) on average with std 0.42.
While both benchmarks include creative images, these results show that Whoops! questions adhere to conventional styles, which do not prompt GPT-4o to stereotype its answers. In contrast, our benchmark’s suggestive questions are intentionally designed to elicit stereotype-averse responses, aligning with our design principles.
D.2.3 Compare to HallusionBench Guan et al. (2023)
HallusionBench consists of 2 parts: Visual Dependent (which focuses on testing a model’s general visual reasoning skills) and Visual Supplement (which evaluates a model’s visual reasoning ability and the balance between parametric memory and image context).
The Visual Supplement part is related to our benchmark, as its questions, like ours, can be answered without visual information. However, the key difference lies in their design. HallusionBench questions rely on parametric memory and strict factual knowledge, e.g., Which country has the most gold medals in the 2024 Olympics?, whereas our questions are based on stereotypes, e.g., A soccer ball is round. This design in HallusionBench significantly constrains the scope of questions and the benchmark’s size, as the official release includes only 50 question pairs. Additionally, it is unclear whether VLM models are expected to rely on their factual knowledge or the inputted, modified chart to answer these questions, as the altered images contradict real-world facts.
Furthermore, all of HallusionBench’s images fall into specific categories: chart, table, map, and OCR, representing a very narrow range of visual information. While the benchmark claims to test a model’s reasoning ability by making subtle changes to these images (e.g., altering a single digit in a chart), reading and understanding such images is already challenging, even without modifications. And these minor changes further complicate the task, and makes the results hard to interpret.
Figure 34: Example of HallusionBench. HallusionBench also has questions which can be answered without images. While it is based on facts instead of stereotypes like ours. Moreover, its images are limited to Chart, Map, OCR and Table.
We show representative images of HallusionBench in this image, which clearly illustrates the very narrow range of image and QA types compared to our benchmark. The modification of the image (from green boxed image to red one) is hard to recognize even for humans.
D.3 Human Study
For human evaluation, we hired Ph.D.-level candidates to participate in testing. They were asked to answer questions with a single image provided each time, and the QIAs were randomly shuffled to avoid any sequential context. To ensure efficiency, we conducted an oral test instead of a written one, recording their responses. After the test, we updated the synonym sets for the QIAs based on their answers. Every test session lasted approximately 1.5 hours, and all the participants expressed confidence in their answers.

### v3

@@
 Abstract
-Despite recent advances in Vision-Language Models (VLMs), many still over-rely on visual language priors existed in their training data rather than true visual reasoning.
-To examine the situation, we introduce ViLP, a visual question answering (VQA) benchmark that pairs each question with three potential answers and three corresponding images: one image whose answer can be inferred from text alone, and two images that demand visual reasoning.
-By leveraging image generative models, we ensure significant variation in texture, shape, conceptual combinations, hallucinated elements, and proverb-based contexts, making our benchmark images distinctly out-of-distribution.
-While humans achieve near-perfect accuracy, modern VLMs falter; for instance, GPT-4 achieves only 66.17% on ViLP.
-To alleviate this, we propose a self-improving framework in which models generate new VQA pairs and images, then apply pixel-level and semantic corruptions to form “good-bad” image pairs for self-training.
-Our training objectives compel VLMs to focus more on actual visual inputs and have demonstrated the effectiveness in enhancing the performance of open-source VLMs, including LLaVA-v1.5 and Cambrian.
+Despite recent advances in Vision-Language Models (VLMs), they may over-rely on visual language priors existing in their training data rather than true visual reasoning.
+To investigate this, we introduce ViLP, a benchmark featuring deliberately out-of-distribution images synthesized via image generation models and out-of-distribution Q&A pairs.
+Each question in ViLP is coupled with three potential answers and three corresponding images: one that can be resolved by text priors alone and two that demand visual reasoning.
+Although, humans achieve near-perfect accuracy, modern VLMs falter; for instance, GPT-4 achieves only 66.17% on ViLP.
+To alleviate this, we propose a self-improving framework in which models generate new VQA data, then apply pixel-level and semantic corruptions to form “good-bad" image pairs for self-training.
+Our training objectives compel VLMs to focus more on the actual visual inputs, and we demonstrate their effectiveness in boosting the performance of open-source VLMs, including LLaVA-v1.5 and Cambrian.
+Machine Learning, ICML
+[…]
 1 Introduction
-Vision-Language Models (VLMs) have advanced text-image interaction, serving as a cornerstone in bridging the gap between visual and textual data (Achiam et al., 2023; Team et al., 2023).
-However, a persistent challenge for learning-based models, such as ResNets and CLIPs, is their reliance on strong priors overly learned in the training data, sometimes overlooking visual cues for answering questions (Agrawal et al., 2016; Prabhu et al., 2023).
-For example, when shown an image of a hexagonal moon (Figure 1), a model might incorrectly assert that “the moon is round" due to strong language priors.
-Similarly, these models may adhere to learned visual priors (Thrush et al., 2022; Sterz et al., 2024), making it difficult for them to comprehend out-of-distribution visual cues, such as a zebra with atypical spot patterns (Figure 1), where humans would easily discern the difference.
-This raises an important question: do today’s VLMs still over-rely on learned visual language priors?
+Vision-Language Models (VLMs) have advanced text-image interaction, bridging the gap between visual and textual data (Achiam et al., 2023; Team et al., 2023).
+However, a persistent challenge for learning-based models, such as ResNets and CLIPs, lies in their reliance on learned priors from the training data, sometimes overlooking visual cues when answering questions (Agrawal et al., 2016; Prabhu et al., 2023).
+For example, when shown a torus-shaped soccer ball (Figure 1), a model might incorrectly identify it as a sphere due to strong language priors.
+Simultaneously, these models may adhere to learned visual priors (Thrush et al., 2022; Sterz et al., 2024), making it difficult to comprehend out-of-distribution visual cues, such as a zebra with atypical spot patterns (Figure 1), which humans would easily discern.
+This raises an important question: do today’s VLMs still over-rely on learned visual language priors, especially given that they rely on far fewer image-text pairs compared to the internet-scale text corpora used for pretraining?
 To investigate this, we probe the Visual Language Priors of VLMs by constructing Question-Image-Answer (QIA) triplets that deliberately deviate from the training data distribution.
-Unlike existing benchmarks that typically rely on internet-sourced images, which inadvertently favor the visual language priors embedded in training data, we utilize modern image generation models such as DALL·E-3 (Ramesh et al., 2021) and Flux to synthesize out-of-distribution images, supporting uncommon answers.
-To further challenge the priors of VLMs, we amplify the language priors in questions by introducing distractor facts, making the questions more suggestive.
+Unlike existing benchmarks that typically rely on internet-sourced images (Goyal et al., 2017; Tong et al., 2024), which inadvertently favor the visual language priors embedded in the training data of VLMs, we utilize modern image generation models, including DALL·E-3 (Ramesh et al., 2021) and Flux, to synthesize out-of-distribution images supporting also out-of-distribution answers.
+Using the image generation models also helps our QIAs exhibit notable variation in texture, shape, conceptual combinations, hallucinated elements, and proverb-based contexts.
 Our benchmark, ViLP, contains 300 carefully designed questions, each paired with three distinct answers: a Prior Answer and two Test Answers, resulting in a total of 900 QIA triplets.
-Each question is structured to present a distractor fact followed by a question.
+To further challenge the priors of VLMs, we amplify language priors in questions by introducing distractor facts: each question is structured to present a distractor fact followed by a question.
 The distractor fact directly leads to the Prior Answer.
 In contrast, the two Test Answers are crafted to challenge the priors by requiring both textual and visual cues for accurate reasoning.
-While human participants can differentiate between all three answers, current VLMs face considerable difficulty, as evidenced by a significant performance drop in our benchmarks, with GPT-4o (OpenAI, 2024) scoring only 66.17%.
-Inspired by ViLP  results, we propose a self-improving approach to encourage VLM reasoning more relying on visual inputs.
-Our method involves self-generating VQAs using image generation and editing models (Podell et al., 2023; Ren et al., 2024; Brooks et al., 2023a) and introducing controlled corruptions to create “good" and “bad" image pairs for DPO-like training (Rafailov et al., 2024).
-We theoretically prove that our proposed objective optimizes an upper bound of the RLHF objective (Ouyang et al., 2022), and experiments on open-sourced VLMs, including LLaVA-v1.5 (Liu et al., 2024) and Cambrian (Tong et al., 2024), demonstrate its effectiveness.
+While human participants achieved $\sim$98accuracy easily, current VLMs exhibit considerable difficulty, as evidenced by a significant performance drop on our benchmarks, with GPT-4o (OpenAI, 2024) scoring only 66.17%.
+Motivated by the results of ViLP, we propose Image-DPO, a self-improving approach to enhance VLM visual reasoning performance by increasing reliance on visual inputs.
+Our method employs self-generated VQAs using image generation and editing models (Podell et al., 2023; Ren et al., 2024; Brooks et al., 2023a) and applies controlled corruptions to create “good" and “bad" image pairs for DPO-like training (Rafailov et al., 2024).
+Experiments with open-source VLMs, including LLaVA-v1.5 (Liu et al., 2024a) and Cambrian (Tong et al., 2024), demonstrate its effectiveness.
+Moreover, we theoretically show that our objective optimizes an upper bound of the RLHF objective (Ouyang et al., 2022).
+The proposed ViLP dataset, Image-DPO code, and our synthetic data appear in https://huggingface.co/ViLP.
 Figure 1: Sample data from ViLP.
@@
 The Prior Answer (first column) can be directly inferred from the question, while Test Answers (second &\&& third column) rely on visual cues.
-All of our answers are designed to be single words, allowing for evaluation without the need for LLMs.
-We have developed a robust synonym and plural detection pipeline to ensure that open-ended responses do not interfere with the evaluation process.
-Please refer to Appendix A for more data samples from ViLP.
+Our answers are designed to be single words, and both the model and human evaluators are tasked with open-domain answering, rather than selecting from predefined options.
+To support this, we have developed a robust synonym and plural detection pipeline, ensuring that open-ended responses do not hinder the evaluation process.
+This approach also enables evaluation without relying on LLMs.
+Please refer to Appendix A.1 for more data samples from ViLP.
+Furthermore, we include both qualitative and quantitative comparison results with Winoground (Thrush et al., 2022), Whoops!(Bitton-Guetta et al., 2023a), and HallusionBench(Guan et al., 2023) in Appendix A.3.
 2 Related Work
-VQA Dataset: There have been long-term efforts to establish Visual Question Answering (VQA) datasets from various perspectives, including general visual question answering  (Agrawal et al., 2015; Gurari et al., 2018; Fu et al., 2023; Liu et al., 2023e; Li et al., 2023a; Yu et al., 2023b; Liu et al., 2024), reading text or charts (Singh et al., 2019a; Mathew et al., 2020; 2021; Masry et al., 2022), complicated reasoning (Lu et al., 2022; 2023), probing compositions (Hudson & Manning, 2019; Ma et al., 2022; Thrush et al., 2022; Hsieh et al., 2023; Li et al., 2024a), studying hallucinations (Rohrbach et al., 2018; Li et al., 2023c; Guan et al., 2023), common sense reasoning (Bitton-Guetta et al., 2023b; a), and more (Majumdar et al., 2024; Sterz et al., 2024).
-This paper proposed a benchmark to evaluate the visual reasoning abilities of VLMs when the questions, answers and images defy common visual language priors.
-Our approach follows the balanced dataset design in Goyal et al.
-(2017), where each question is associated with three answers: the first aligns with language priors, while the latter two deviate from them, requiring reliance on visual cues for correct answering.
-By leveraging state-of-the-art image generation models, our benchmark creates images that defy priors much more effectively than previous works using internet images, such as Goyal et al.
-(2017); Tong et al.
-(2024).
-Moreover, unlike the “trick” category in Sterz et al.
-(2024), our process first generates question-answer pairs and then synthesizes images that meet the specified conditions, resulting in more challenging visual reasoning tasks.
-More comparisons with existing datasets are included in Appendix D.2.
-Vision Language Models and Language Priors: Multimodal reasoning is essential for machine intelligence, with vision-language models (VLMs) integrating visual perception, text reasoning, instruction following, and generation for complex tasks (Tan & Bansal, 2019; Li et al., 2019; Kim et al., 2021; Wang et al., 2021b; a; Alayrac et al., 2022; Li et al., 2023b; Chen et al., 2022; Jia et al., 2021; Shen et al., 2021; Singh et al., 2021; Liu et al., 2023c; a; Zhao et al., 2023; Chen et al., 2023; Zhu et al., 2024; Li et al., 2024c; Dai et al., 2023; Li et al., 2024c; Yu et al., 2024; Dai et al., 2024; Deitke et al., 2024).
-Building on the success of language models (Brown et al., 2020; OpenAI, 2023a; Touvron et al., 2023a; b; Chiang et al., 2023) and pre-trained visual encoders (Radford et al., 2021; Desai & Johnson, 2020; Caron et al., 2021; Chen et al., 2024), many recent advancements leverage vision-language paired datasets (Liu et al., 2024; Tong et al., 2024) to fine-tune intermediate connectors between LLMs and visual encoders (Liu et al., 2024).
-However, these paired datasets are significantly smaller than the large text corpora used for LLM pre-training (OpenAI, 2023b; Soldaini et al., 2024), and many approaches do not update the visual encoder and LLM parameters during fine-tuning, which can lead to persistent language biases.
-Specifically, the outputs of VLMs become dominated by the language priors learned from the pre-training corpus, reducing the importance of visual inputs (Thrush et al., 2022; Sterz et al., 2024).
-This issue becomes even more critical when VLMs are presented with deliberately generated images that challenge these biases, as demonstrated in our work.
-Previous methods and datasets (Goyal et al., 2016; Agrawal et al., 2017; Dancette et al., 2021; Wu et al., 2022; Ramakrishnan et al., 2018; Gouthaman & Mittal, 2020) have attempted to assess these issues by studying visual backbones and curating datasets using simulators (Johnson et al., 2016) or internet images (Zhang et al., 2016).
-This paper examines these challenges by creating a novel VQA benchmark, featuring carefully designed questions, fact-based distractors, and rare-distribution answers.
-Furthermore, we utilize advanced image generation techniques to produce realistic images that subvert common language biases as shown in Figure 1.
-Self-Rewarding VLM: Self-rewarding LLM (Yuan et al., 2024) has demonstrated the potential of using LLMs to generate and optimize data through directed reference optimization (Rafailov et al., 2024), leading to self-improvements.
-This line of research has been extended to VLMs, where new answers are generated and rated for DPO training (Zhou et al., 2024a; Deng et al., 2024; Zhou et al., 2024c; Wang et al., 2024b; a).
-Our findings resonate with this body of work on self-rewarding VLMs; however, we present two key distinctions: (1) our proposed image-DPO method generates multiple images for a single question-image pair, rather than generating multiple answers and associated ratings; (2) unlike previous methods that focused solely on existing images, both our proposed image-DPO and text-DPO pipelines produce a diverse array of new images with pre-trained image generative models, including SDXL (Podell et al., 2023), GroundedSAM (Ren et al., 2024), and InstructPix2Pix (Brooks et al., 2023a).
-Specifically, as illustrated in Figure 6, our proposed text-DPO is similar to the self-rewarding VLMs mentioned earlier but incorporates the generation of new images using pre-trained image generation models.
-In contrast, our image-DPO deliberately corrupts images to create multiple degraded versions, which are used as rejected data during DPO training.
-This is intended to compel the VLM to rely more on the visual input for accurate generation.
+VQA Dataset: Significant efforts have produced VQA datasets from various angles, including general VQA (Agrawal et al., 2015; Gurari et al., 2018; Fu et al., 2023; Liu et al., 2023d; Li et al., 2023a; Yu et al., 2023b; Liu et al., 2024a), reading text or charts (Singh et al., 2019a; Mathew et al., 2020, 2021; Masry et al., 2022), complex reasoning (Lu et al., 2022, 2023), composition probing (Hudson & Manning, 2019; Ma et al., 2022; Thrush et al., 2022; Hsieh et al., 2023; Li et al., 2024a), hallucinations (Rohrbach et al., 2018; Li et al., 2023c; Guan et al., 2023), common-sense reasoning (Bitton-Guetta et al., 2023b, a), and more (Majumdar et al., 2024; Sterz et al., 2024).
+We propose a benchmark that tests VLMs’ visual reasoning when questions, answers, and images defy common patterns.
+Following the balanced dataset design of (Goyal et al., 2017), each question is accompanied by three answers: one aligns with language priors, and two deviate, prompting reliance on visual cues.
+By leveraging state-of-the-art image generation models, our benchmark challenges these priors more effectively than previous datasets built from internet images (Goyal et al., 2017; Tong et al., 2024).
+Furthermore, unlike the “trick” category in (Sterz et al., 2024), we first generate question-answer pairs before synthesizing images under specified constraints, creating more difficult visual reasoning tasks.
+Comprehensive comparisons with existing datasets appear in Appendix A.3.
+Vision Language Models and Language Priors: Multimodal reasoning is crucial for machine intelligence, with VLMs integrating visual perception, text reasoning, instruction following, and generation for complex tasks (Tan & Bansal, 2019; Li et al., 2019; Kim et al., 2021; Wang et al., 2021b, a; Alayrac et al., 2022; Li et al., 2023b; Chen et al., 2022; Jia et al., 2021; Shen et al., 2021; Singh et al., 2021; Liu et al., 2023c, a; Zhao et al., 2023a; Chen et al., 2023; Zhu et al., 2024; Li et al., 2024c; Dai et al., 2023; Li et al., 2024c; Yu et al., 2024; Dai et al., 2024; Deitke et al., 2024).
+Inspired by the success of large language models (Brown et al., 2020; OpenAI, 2023a; Touvron et al., 2023a, b; Chiang et al., 2023) and pre-trained visual encoders (Radford et al., 2021; Desai & Johnson, 2020; Caron et al., 2021; Chen et al., 2024), many recent methods leverage relatively small vision-language paired datasets (Liu et al., 2024a; Tong et al., 2024) to fine-tune connectors between LLMs and visual backbones (Liu et al., 2024a).
+However, these datasets are far smaller than the vast text corpora for LLM pre-training (OpenAI, 2023b; Soldaini et al., 2024), and freezing the visual encoder and LLM parameters often preserves language biases, causing visual inputs to be overshadowed (Thrush et al., 2022; Sterz et al., 2024).
+This challenge is amplified by deliberately generated images that expose such biases, as shown in our study.
+Previous works and datasets (Goyal et al., 2016; Agrawal et al., 2017; Dancette et al., 2021; Wu et al., 2022; Ramakrishnan et al., 2018; Gouthaman & Mittal, 2020) addressed these issues with curated simulators (Johnson et al., 2016) or internet imagery (Zhang et al., 2016).
+In this paper, we present a novel VQA benchmark featuring carefully designed questions, fact-based distractors, rare-distribution answers, and image generation techniques to produce realistic visuals that challenge learned visual language priors (Figure 1).
+Self-Rewarding VLM: Self-rewarding LLM (Yuan et al., 2024) has shown that LLMs can generate and improving themselves in the process via Directed Peference Optimization (DPO) (Rafailov et al., 2024).
+This approach extends to VLMs by generating new answers for DPO training (Zhou et al., 2024a; Deng et al., 2024; Zhou et al., 2024b; Wang et al., 2024c, a; Yue et al., 2024; Liu et al., 2024b; Xiao et al., 2024).
+Our work aligns with these self-rewarding VLMs but differs in two key ways: (1) our proposed Image-DPO generates multiple images for a single question-image pair (rather than multiple answers); (2) rather than relying solely on existing images, Image-DPO pipeline creates diverse new images using pre-trained generative models (SDXL (Podell et al., 2023), GroundedSAM (Ren et al., 2024), InstructPix2Pix (Brooks et al., 2023a)).
+Furthermore, Image-DPO deliberately corrupts images to produce multiple degraded versions that serve as rejected data in DPO training.
+Concurrent works (Wang et al., 2024a; Xie et al., 2024) explore similar methodologies but lack a benchmark to verify enhanced visual focus, fail to establish theoretical connections between their proposed objective and DPO, and utilize limited image transformations (e.g., only randomly cropping).
+In contrast, we introduce ViLP (Section 3) to assess visual reasoning and provide theoretical foundations (Appendix B), alongside multi-category image corruptions (semantic editing, Gaussian blurring, pixelation).
 3 ViLP  Benchmark
 3.1 Design Principles
-"What’s the tall animal with a long neck?" Even without any visual cues, humans readily infer that the answer is "giraffe", leveraging language context and commonsense knowledge.
-But is it necessarily a giraffe?
-As shown in the bottom-right of Figure 1, there may be an elephant or a camel with a very long neck and much taller than giraffes.
-This example highlights a potential issue with current large Vision-Language Models (VLMs).
-As VLMs are built upon Large Language Models (LLMs) trained on web-scale corpora and are fine-tuned with comparatively much smaller image-text datasets, they may develop a strong bias toward certain visual language priors (especially language) presented in training data and overlook real visual input (Thrush et al., 2022; Sterz et al., 2024).
-Regarding visual language priors, we focus on strong language priors that confuse VLMs to derive direct answers solely from textual cues.
-Simultaneously, we aim to create images that are distinctly out-of-distribution to challenge potential visual priors which make the model overlook the critical visual cues such as the unusual spots patterns in the Zebra shown in Figure 1.
-To assess the visual reasoning abilities of VLMs against visual language priors, we introduce a new benchmark, ViLP, featuring carefully constructed out-of-distribution question-image-answer (QIA) triplets.
-These triplets are designed based on the below principles ensuring that VLMs rely on both textual and visual cues, rather than learned priors, to arrive at correct answers.
-  • Text-Only Inference: Create questions that, when considered alone, lead to a high-confidence answer based solely on textual information.
-  • Visual Influence: When an accompanying image is present, the correct answer differs significantly from the text-only response, sometimes contradicting common sense, ensuring the critical role of visual context in ensuring a high-confidence result.
-Mathematically, with $Q$ denoting a question, $I$ denoting an image, $A=\{a_{\text{prior}},\cdots,a_{\text{test}},\cdots\}$ being the set of all possible answers, we use $P(a\mid Q)$ and $P(a\mid Q,I)$ to denote the probability of answer $a$ given the question $Q$ alone, and the probability of answer $a$ given both the question $Q$ and the image $I$, respectively.
-Here $p$ is a prior model, which could either be human’s cognition prior ($P_{\text{human}}$) or a VLM/LLM’s learned visual language prior ($P_{\theta}$).
-For constructing our benchmark, we used the following objectives:
-Criteria One: Without considering the image, the question $Q$ should lead to a high-confidence answer $a_{\text{prior}}$, where $\delta_{1}$ is a high-confidence threshold.
-$a_{\text{prior}}$ usually satisfies common knowledge, such as “the moon is round" and “Leonardo da Vinci painted the Mona Lisa" as shown in Figure 1.
+“What’s the tall animal with the longest neck shown?” Humans readily guess “giraffe” based on learned priors, , yet as shown in the bottom-right of Figure 1, it could be an elephant or camel – where visual reasoning corrects the answer.
+This highlights a potential shortfall in Vision-Language Models (VLMs), which may over-rely on learned visual language priors instead of true visual reasoning, particularly since VLMs are typically fine-tuned on limited image-text data, which is several orders of magnitude smaller than the trained text corpus (Liu et al., 2024a; Tong et al., 2024).
+Specifically for the scope of visual language priors in this paper, we target (1) strong language priors that lead VLMs to derive answers solely from text, and (2) potential visual priors causing models to overlook critical uncommon visual cues (e.g., unusual zebra spots in Figure 1).
+To evaluate how VLMs handle learned visual language priors, we introduce ViLP, a specialized benchmark of out-of-distribution Question-Image-Answer (QIA) triplets guided by two core principles.
+First, Text-Only Inference ensures that each question can be answered with high confidence using textual clues alone.
+Second, Visual Influence requires that the correct answer—sometimes contradicting common sense—only emerges once an out-of-distribution image is considered.
+By forcing models to integrate both textual and visual information, ViLP reveals whether they truly engage in visual reasoning or merely rely on memorized patterns.
+Mathematically, let $Q$ be a question, $I$ an image, and $A={a_{\text{prior}},\dots,a_{\text{test}},\dots}$ the set of possible answers.
+We define $P(a\mid Q)$ as the probability of answer $a$ given $Q$ alone and $P(a\mid Q,I)$ as the probability given both $Q$ and $I$.
+We consider a prior model $p$, which may represent either human cognition ($P_{\text{human}}$) or a VLM/LLM’s learned visual-language prior ($P_{\theta}$).
+For constructing our benchmark, we used the following guidances:
+Criterion One: The question $Q$ alone should strongly favor $a_{\text{prior}}$, where $\delta_{1}$ is a high-confidence threshold.
+$a_{\text{prior}}$ usually satisfies common knowledge, such as “soccer ball is a sphere" and “Einstein created $E=mc^{2}$" (Figure 1).
 $P(a_{\text{prior}}\mid Q)\geq\delta_{1}$ (1)
-Criteria Two: With the inclusion of the image $I$, the correct answer becomes $a_{\text{test}}$, where $\delta_{2}$ is another high-confidence threshold.
-Also, $a_{\text{test}}$ is significantly different from $a_{\text{prior}}$ to ensure that the image has a substantial impact on the answer, where $D$ is a divergence measure (e.g., Kullback-Leibler divergence), and $\delta_{3}$ is a threshold indicating significant difference.
-For instance, the image in the second-to-last row and column of Figure 1 clearly shows the moon as a hexagon.
-$P(a_{\text{test}}\mid Q,I)\geq\delta_{2}\quad\text{and}\quad D\big{(}P(a_{\text{prior}}\mid Q),P(a_{\text{test}}\mid Q,I)\big{)}\geq\delta_{3}$ (2)
-Criteria Three: We design answer $a_{\text{test}}$ to be a rare choice, one that is unlikely to be inferred from the question $Q$ alone.
-In contrast, answer $a_{\text{prior}}$ should be clearly incorrect when considering both the question $Q$ and the image $I$, where $\delta_{4}$ represents a low-confidence threshold.
-Typically, $a_{\text{test}}$ will be out of context.
-For example, in the image from the 3rd row and 2nd column of Figure 1, Vincent van Gogh serves as $a_{\text{test}}$, inferred from the image, while Leonardo becomes contradictory after incorporating the additional image condition.
-$P(a_{\text{test}}\mid Q)\leq\delta_{4}\quad\text{and}\quad P(a_{\text{prior}}\mid Q,I)\leq\delta_{4}$ (3)
+Criterion Two: With the image $I$, the correct answer shifts to $a_{\text{test}}$, where $\delta_{2}$ is another high-confidence threshold.
+Also, $a_{\text{test}}$ is significantly different from $a_{\text{prior}}$ to ensure that the image has a substantial impact on the answer, where $D$ is a divergence measure, and $\delta_{3}$ is a threshold indicating significant difference.
+For instance, the image in the 1st row and the 3rd column of Figure 1 turns the answer to torus.
+$P(a_{\text{test}}\mid Q,I)\geq\delta_{2},\,D\big{(}P(a_{\text{prior}}\mid Q),P(a_{\text{test}}\mid Q,I)\big{)}\geq\delta_{3}$ (2)
+Criterion Three: The answer $a_{\text{test}}$ should be rare and unlikely from $Q$ alone, while $a_{\text{prior}}$ becomes clearly incorrect when considering $I$.
+This is enforced by a low-confidence threshold $\delta_{4}$ (e.g., Newton as $a_{\text{test}}$ inferred from the image, contradicting Einstein, shown by the image from the 3rd row and 2nd column of Figure 1).
+$P(a_{\text{test}}\mid Q)\leq\delta_{4},\,P(a_{\text{prior}}\mid Q,I)\leq\delta_{4}$ (3)
 In constructing ViLP, we use human cognition prior $P_{\text{human}}$ as guidance for its design.
 We evaluated the behavior of the learned priors of VLMs/LLMs by benchmarking $P_{\theta}$ against the human prior $P_{\text{human}}$.
-This benchmarking process implicitly measuring the divergence between $P_{\text{human}}$ and $P_{\theta}$.
 3.2 Question-Image-Answer Generation
-As outlined in Criteria Three, we aim for the test answer, $a_{\text{test}}$, to appear highly improbable when considered solely in the context of the question $Q$, yet become a clear and confident choice when paired with the image $I$.
-This objective presents a significant challenge in finding images $I$, as such images may not follow common knowledge and not naturally exist in the real world or the internet.
-Recent advancements in generative models offer a solution.
-By utilizing advanced image generation models such as DALL·E-3 (Ramesh et al., 2021) and Flux, which follow text prompts to synthesize images, we can create images that blend unusual elements.
-This capability allows us to generate images with the necessary visual coherence to satisfy the specific demands of the question and answer, where we create in a way to disobey language priors.
-To further ensure the generation of question-image-answers (QIAs) that align with the principles outlined in Section 3.1, our benchmark dataset design incorporates substantial human effort, assisted by state-of-the-art language models such as OpenAI-o1 and Claude-3.5-Sonnet, with additional details provided in Appendix D.1.
-Briefly, for each question, we designed three different answers where the first answer $a_{\text{prior}}$ could be inferred solely from the question text.
-The other two answers were specifically designed against language priors, challenging the model to interpret visual cues despite strong textual priors.
-During evaluation, we focus on the model’s performance on these second and third answers, i.e., $a_{\text{test}}$.
-After crafting the initial questions and answers, we leverage GPT-4 to generate text prompts that guide the creation of diverse images at scale.
-This facilitates the production of varied visual data for the next stage, where high-quality QIA triplets are selected through human review and filtering.
-Both the text and images are refined as necessary when misalignment or ambiguity are identified.
-We ensure that the final QIA triplets are clear and easily interpretable by humans, as evidenced by the results of our human evaluation results in Table 2.
-This iterative process involves several cycles of adjustments to achieve the desired quality.
-Throughout the process, we encountered two main challenges: (1) generating diverse, out-of-distribution, while reasonable question-answer pairs are difficult, as models like OpenAI-o1 and Claude-3.5-Sonnet tend to converge on a limited set of responses, requiring significant human input to introduce variety; and (2) generating images that against with specific priors using state-of-the-art image generation models is also challenging.
-In some cases, it was necessary to generate hundreds of images to find just one that accurately matched the question and answer, while avoiding ambiguity.
-Table 1: Categorical information.
+Following Criterion Three, $a_{\text{test}}$ should be highly improbable based on $Q$ alone yet the correct choice when paired with $I$.
+Since such images do not exist naturally, we use generative models like DALL·E-3 (Ramesh et al., 2021) and Flux to blend unusual elements that override typical language priors.
+We incorporate substantial human input and leverage advanced LLMs such as OpenAI-o1 and Claude-3.5-Sonnet to ensure alignment with all the criteria.
+More details are provided in Appendix A.2.
+(a) Image-DPO vs.
+Text-DPO: In the green dashed box, we illustrate Image-DPO, which uses a single Q&A pair paired with multiple corrupted images.
+In contrast, the purple dashed box presents Text-DPO, involving a single input image paired with multiple, distinct Q&A pairs.
+(b) Illustration of Image DPO.
+We construct chosen and rejected pairs by corrupting the image with a set of perturbations while keeping the Q&A unchanged.
+Perturbations include semantic editing, Gaussian blurring, and pixelation.
+The mathematical formulations and implementation details are provided in Appendix B and Appendix C, respectively.
+For each question, we design three answers: one $a_{\text{prior}}$ inferred solely from $Q$, and two $a_{\text{test}}$ that defy language priors, requiring visual cues for correctness.
+We rely on GPT-4 to generate text prompts, produce large-scale images, and then conduct human filtering and refinement.
+This process faces two main challenges: (1) producing diverse out-of-distribution QA pairs, and (2) synthesizing images that defy specific priors, sometimes necessitating hundreds of samples to find one that accurately matches $Q$ and $a_{\text{test}}$.
+Table 1: Category Data.
 Type Frequency
@@
 Proverbs 17
-As a result, we created a total of 300 questions, each paired with three distinct image-answer combinations, resulting in 900 IQA sets.
-These questions span a broad range of topics, from low-level visual recognition tasks such as texture and shape, to higher-level visual reasoning involving conceptual combinations, hallucinated components, and proverbs.
-To reinforce the text priors, we also provide a distractor fact before the question, offering additional context for the prior answer $a_{\text{prior}}$ (the one inferred without the image).
-Some categorical information is shown on the right Table 1.
-Most of our data are not confined to a single category; instead, they frequently combine multiple categories to challenge even the most advanced VLMs.
-For instance, in the moon example (6th-row, 2nd-col), we explore shapes and conceptual combinations, such as a round moon paired with a hexagon versus a hexagonal moon.
-Similarly, the umbrella example (7th-row, 1st-col) incorporates both conceptual combinations and hallucinated components, blending elements like a plunger and a blender.
-On average, our data span 1.6 categories per question.
+Ultimately, we curated 300 questions, each paired with three distinct image-answer sets, totaling 900 QIA triplets.
+These cover a broad range from low-level recognition (texture, shape) to high-level reasoning (conceptual combinations, hallucinated components, proverbs).
+Table 1 summarizes their categorical distribution, with each question spanning an average of 1.6 categories.
+To reinforce text priors, we present a distractor fact before the question.
+Rigorous human review ensures that all final QIA triplets are clear and interpretable, as reflected by our human evaluation results in Table 2.
 3.3 Dataset Evaluation
@@
 4 Image DPO
-Inspired by our benchmark results, we propose Image DPO, a self-improvement method for enhancing VLMs’ visual reasoning, featuring a new objective and a data generation pipeline using VLMs themselves and pre-trained image models.
+Inspired by our benchmark, we propose Image DPO, a self-improvement method for enhancing VLMs’ visual reasoning, featuring a new objective and a data generation pipeline using VLMs themselves and pre-trained image models.
 4.1 Objective
-Figure 2: Illustration of Image DPO.
-We construct chosen and rejected pairs by corrupting the image with a set of perturbations while keeping the text (questions and answers) unchanged.
-This setup encourages that the model more focus on the image to differentiate between the chosen and rejected pairs, as both pairs contain identical textual information.
-The perturbations applied to the images include semantic editing, Gaussian blurring, and pixelization.
-More details in Appendix C.2.
-Existing approaches for VLM self-improvement usually follow the Direct Preference Optimization (DPO) (Rafailov et al., 2024) used in Large Language Models (LLMs), where the model is trained to distinguish between good and bad answers for a fixed image and question.
-However, this straightforward adaptation is not vision-centered, as the model can still distinguish good and bad answers from the text alone without needing to analyze the image.
-In contrast, we propose Image DPO, a vision-focused objective that creates good and bad question-image-answer pairs by corrupting the image while keeping the question and answer unchanged (Figure 2).
-Mathematically, given the image $I_{w}$, question $Q$ and the corresponding answer $A$, we generate a corrupted image $I_{l}$ by various image editing operations including gaussian blur, pixelation or semantic editing.
-With the corrupted image $I_{l}$, the triplet of $Q$, $I_{l}$, and $A$ should form a bad question-image-answer pair compared to the original triplet of $Q$, $I_{w}$, and $A$.
-We train our model to distinguish between good and bad question-image-answer triplets using the objective function:
-$L(\theta,\theta_{\text{ref}})=-\mathbb{E}_{Q,I_{w},I_{w},A\sim S}\left[\log\sigma(\alpha\frac{\pi_{\theta}(A\mid Q,I_{w})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{w})}-\alpha\frac{\pi_{\theta}(A\mid Q,I_{l})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{l})})\right]$ (4)
-Here, $\theta$ represents the VLM model being trained, $\theta_{\text{ref}}$ is the reference VLM model (typically the previous version of $\theta$), $S$ is the training data consisting of good and bad question-image-answer triplets, $\sigma$ is the sigmoid function, and $\alpha$ is a positive constant.
-In Appendix B, we demonstrate that this objective optimizes an upper bound of the RLHF objective (Ouyang et al., 2022), sets it apart from the original DPO objective.
-Intuitively, since the textual inputs and outputs are identical in both good and bad cases, the gradients of this objective push the model to rely more on the vision branch, driving a shift in gradient direction when processing high-quality images $I_{w}$ compared to corrupted images $I_{l}$, as illustrated in Appendix Figure 9.
-This behavior encourages the model to focus more on image inputs rather than relying solely on text-based reasoning, thereby enhancing its performance on visually-related tasks.
+Figure 3: Qualitative examples. We show the results from GPT-4o, Claude-3.5-Sonnet, Gemini-1.5-Pro, and Llama-3.2-Vision-90B for some challenging cases.
+Existing approaches for VLM self-improvement follow the way used in DPO paper (Rafailov et al., 2024), where the model is trained to distinguish between good and bad answers for a fixed image and question (Figure 2(a) right).
+However, this straightforward adaptation may not the best for vision models, as the model sometimes distinguish good and bad answers from the text alone without needing to analyze the image.
+In contrast, we propose Image DPO, a vision-focused objective that creates good and bad question-image-answer pairs by corrupting the image while keeping the question and answer unchanged (Figure 2(a) left).
+An example of our synthetic data is illustrated in Figure 2(b).
+Formally, given an image $I_{w}$, a question $Q$, and its corresponding answer $A$, we generate a corrupted image $I_{l}$ via image-editing operations, including Gaussian blur, pixelation, or semantic modifications.
+The triplet $(Q,I_{l},A)$ forms a degraded question-image-answer pair compared to $(Q,I_{w},A)$.
+We train the model to distinguish between good and bad triplets using the objective below, where $\theta$ is the target VLM, $\theta_{\text{ref}}$ is the reference VLM (typically an earlier version of $\theta$), $S$ is the dataset of good and bad triplets, $\sigma$ is the sigmoid function, and $\alpha$ is a scaling factor.
+$L(\theta,\theta_{\text{ref}})$ $=-\mathbb{E}_{Q,I_{w},I_{l},A\sim S}\Bigl{[}$ (4)
+$\quad\log\sigma\Bigl{(}\alpha\frac{\pi_{\theta}(A\mid Q,I_{w})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{w})}\;-\;\alpha\frac{\pi_{\theta}(A\mid Q,I_{l})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{l})}\Bigr{)}\Bigr{]}$
+Intuitively, since the textual inputs and outputs are identical in both good and bad cases, the gradients of this objective push the model to rely more on the vision branch, driving a shift in gradient direction when processing normal images $I_{w}$ compared to corrupted images $I_{l}$ (Figure 13).
+This behavior encourages the model to focus more on image inputs rather than relying solely on text-based reasoning, thereby enhancing its performance on visual-related tasks.
+Our experiments demonstrate Image-DPO objective (Eq.
+4) outperforms varies self-improve VLM baselines on ViLP.
+Proposition 1.
+Let $\mathcal{L}_{\mathrm{RLHF}}(\pi_{\theta},\pi_{\mathrm{ref}};\mathcal{S})$ be the KL-constrained reward maximization objective in Appendix Eq.
+7 (Rafailov et al., 2024), where the dataset $\mathcal{S}=\{(Q,A,I_{w},I_{l})\}$ contains good images $I_{w}$ and corrupted images $I_{l}$.
+Let $\mathcal{L}_{\mathrm{ImageDPO}}(\pi_{\theta},\pi_{\mathrm{ref}};\mathcal{S})$ be the objective from Eq.
+4, which compares $\bigl{(}Q,I_{w},A\bigr{)}$ against $\bigl{(}Q,I_{l},A\bigr{)}$.
+Then for any policy $\pi_{\theta}$ and reference model $\pi_{\mathrm{ref}}$, we have
+$\mathcal{L}_{\mathrm{RLHF}}(\pi_{\theta},\pi_{\mathrm{ref}};\mathcal{S})\;\;\leq\;\;\mathcal{L}_{\mathrm{ImageDPO}}(\pi_{\theta},\pi_{\mathrm{ref}};\mathcal{S}).$
+Proof Sketch.
+Following (Rafailov et al., 2024), we express the optimal KL-constrained policy in terms of a latent reward function.
+Applying a Bradley–Terry preference model to question-image-answer triplets $\bigl{(}Q,I_{w}/I_{l},A\bigr{)}$ and using Jensen’s inequality yields an upper bound whose minimization is equivalent to $\mathcal{L}_{\mathrm{ImageDPO}}(\pi_{\theta},\pi_{\mathrm{ref}};\mathcal{S})$.
+A full derivation appears in Appendix B.
 4.2 Data Generation
-Training VLMs requires a large amount of question-image-answer (QIA) triplets, which are often scarce and expensive to collect.
-In response, we present a scalable data generation pipeline that creates new QIAs from existing datasets by leveraging VLMs and pre-trained image generation and editing models, as illustrated in Appendix Figure 10.
-Our pipeline is akin to the tool usage of VLMs/LLMs as prior works (Gupta & Kembhavi, 2022; Parisi et al., 2022; Yao et al., 2022; Sur’is et al., 2023), where pre-trained image models are invoked as callable functions.
-Given a seed image, VLMs are tasked with simultaneously selecting appropriate functions (e.g., image generation or editing models) and generating corresponding instructions.
+Training VLMs demands large-scale question-image-answer (QIA) triplets, which are often scarce.
+To address this, we introduce a scalable data generation pipeline (Appendix Figure 14) that repurposes existing image datasets via VLM themselves and image generative models.
+Given a seed image from COCO (Lin et al., 2014), Text2VQA (Singh et al., 2019b), or Visual Genome (Krishna et al., 2017), VLMs are tasked with simultaneously selecting appropriate functions (e.g., image generation or editing models) and generating corresponding instructions.
 These instructions are then used to produce new images, in addition to the seed image, as illustrated in Figure 15.
 The same VLMs are then employed to generate QA pairs for these newly created images.
-Following, we apply the mentioned three types of image corruptions to the generated images, constructing good-bad QIA pairs (i.e., $I_{w}$ and $I_{l}$).
+Following, we apply the mentioned three types of image corruptions to the generated images, constructing good bad pars $(Q,I_{w},A)$ and $(Q,I_{l},A)$.
 Specifically, we employ Stable Diffusion XL (Podell et al., 2023; Rombach et al., 2022) for image generation, and use Instruct-Pix2Pix (Brooks et al., 2023a), and Grounded-SAM (Rombach et al., 2022; Ren et al., 2024) for image editing.
-All of our seed images are sourced from existing datasets, including COCO (Lin et al., 2014), Text2VQA (Singh et al., 2019b), and Visual Genome (Krishna et al., 2017).
-Comprehensive details are provided in Appendix C.
+Example generated data, prompts, and more details are included in Appendix C.
 5 Experiments
-Figure 3: Qualitative examples. We show the results from GPT-4o, Claude-3.5-Sonnet, Gemini-1.5-Pro, and Llama-3.2-Vision-90B for some challenging cases.
-We introduce ViLP, a benchmark designed to evaluate VLMs’ visual reasoning capabilities in the presence of strong visual language priors, consisting of 300 carefully crafted questions.
-Each question is paired with three distinct images and corresponding answers, forming one QIAprior (QIA: question-image-answer) and two QIAtest, resulting in a total of 900 unique QIAs.
-The QIAprior aligns with common knowledge and language priors, making it easy to infer even from text alone.
-The QIAtest, on the other hand, challenges the VLM’s visual reasoning ability against these priors.
-ViLP  includes two evaluation settings: ViLPF (with distractor facts provided) and ViLPP (pure questions without distractor facts).
-For each setting, we report the average accuracy on QIAtest as our focused benchmark Score, and the accuracy on QIAprior as Prior.
-Results are on the Table 2.
-Is the QIA easy for humans?
-We first evaluate our benchmark using human study.
-Humans achieved nearly 100% accuracy on QIAprior and over 98% accuracy on QIAtest, demonstrating that our question-image-answer combinations are non-ambiguous for humans to answer.
-Despite QIAtest being designed as out-of-distribution examples, humans were able to distinguish them.
-Humans performed slightly better on ViLPF-QIAprior when distractor facts were provided, as they could easily identify that these facts aligned with the correct answers.
-Moreover, Score on ViLPF-QIAtest was marginally lower when facts were introduced, as the distractor facts added some noise and caused minor confusion, although the impact of this noise is relatively small.
-These findings are consistent with the design principles of our benchmark.
 Table 2: Benchmarking on ViLP. Please refer to the left text for symbol definitions. †  indicates the model often fails to follow the instructions.
@@
 LLaVA-OneVision-72B 1.67 3.00 5.22 11.67
-Figure 4: Before and after removing distactor facts.
+We introduce ViLP,  a benchmark of 300 questions, each with three unique images and corresponding answers (QIAprior and two QIAtest), yielding 900 QIAs.
+QIAprior aligns with common language priors (easily inferred from text alone), while QIAtest challenges these priors, requiring visual reasoning.
+ViLP features two evaluation settings: ViLPF (with distractor facts) and ViLPP (pure questions).
+We report average accuracy on QIAtest (Score) and on QIAprior (Prior) in Table 2.
+Figure 4: Comparison of benchmark scores under different image transformations. Solid line and dotted line refer to ViLPF-Score and ViLPF-Prior, respectively.
+Is the QIA easy for humans?
+We begin by evaluating our benchmark through a human study.
+Participants achieved nearly perfect accuracy on ViLPF-Prior and over 98% on ViLPF-Score and ViLPP-Score, confirming that our question-image-answer combinations are unambiguous for human interpretation.
+Notably, despite QIAtest being designed as out-of-distribution examples, humans were still able to correctly distinguish them
+Humans performed slightly better on ViLPF-Prior when distractor facts were provided, as they could easily identify that these facts aligned with the correct answers.
+Moreover, ViLPF-Score was marginally lower when facts were introduced, as the distractor facts added some noise and caused minor confusion, although the impact of this noise is relatively small.
+These findings are consistent with the design principles of our benchmark.
+Figure 5: Qualitative results before and after removing distactor facts. GPT-4o and LLaVA-1.5-13B models yield completely opposite behaviors.
 Are our QIAs aligned with the learned priors of VLMs?
-We evaluated GPT-4o’s performance on our questions without providing any images, i.e., GPT-4o (text only) in Table 2.
-Additionally, we eliminate image-related phrases from the prompt during this text-only inference, such as removing expressions like “as shown in the image".
-Remarkably, even without any visual content, the model was able to correctly answer 92.33% of QIAprior when distractor facts were included, demonstrating that our questions are highly suggestive for VLMs, allowing them to answer even without visual input.
-Once removing the distractor facts, the performance decrease to 71.33%, as these facts implicitly pointed to the correct answers in the fact statements.
-It is noteworthy that some powerful models like Claude-3-Opus and several open-source VLMs fail to outperform GPT-4o (text-only) performance on QIAprior, even when provided with images.
-For QIAtest, the success rate of GPT-4o (text-only) dropped dramatically to near zero (0% & 0.17%), indicating that QIAtest are far outside the distribution typically encountered by VLMs or LLMs.
-These observations further validate the design principles of our benchmark: QIAprior can be easily inferred using text alone, while QIAtest is specifically designed to challenge language-based reasoning, requiring reliance on visual cues to arrive at correct answers.
+We tested GPT-4o (text only) on our questions (removing all image references).
+Despite no visual content, it correctly answered 92.33% on ViLPF-Prior.
+The accuracy drops to 71.33% once distractor facts are removed, showing that these facts significantly guide the answer.
+For QIAtest, GPT-4o (text only) accuracy nearly falls to 0% (ViLPF-Score & ViLPP-Score), indicating the QIAtest cannot be answered using text alone.
 How do VLMs perform on our benchmark?
 Although our benchmark questions are distinguishable for humans, they are challenging for VLMs.
-Even the most advanced VLM models like GPT-4o, have a clear performance gap (more than 25%) compared to humans’ performance on QIAtest (Score in ViLPF), indicating the difficulty of these questions for VLMs.
-GPT-4o achieves above 91% accuracy on QIAprior (Prior), which is quite close to human performance.
-While its Score on QIAtest is still low (66.17% versus 98.33%), demonstrating the difficulty of these questions for VLMs.
-We also provide some qualitative examples in Figure 3, showcasing results from both leading commercial and open-source models, including GPT-4o, Claude-3.5-Sonnet, Gemini-1.5-Pro, and Llama-3.2-Vision-90B.
-They encounter challenges when tackling these cases of our ViLP.
-Notably, it is encouraging to see that some open-source models achieved over 60% accuracy on ViLPF-QIAtest, with performance nearing that of their commercial counterparts, including Llama-3.2-Vison and Molmo-72B.
+Even the advanced VLM models like GPT-4o, have a clear performance gap (66.17% v.s.
+98.33%) compared to humans’ performance on ViLPF-Score, indicating the difficulty of these questions for VLMs.
+Claude-3.5-Sonnet achieved the best score 70%, while most of the commercial VLMs are below 60%.
+Figure 3 highlights sample outputs from top commercial and open-source models, including GPT-4o, Claude-3.5-Sonnet, Gemini-1.5-Pro, and Llama-3.2-Vision-90B.
+They face significant challenges when addressing these cases in our ViLP, whereas humans can arrive at correct answers after consideration.
+Notably, it is encouraging to see that some open-source models achieved over 60% accuracy on ViLPF-Score, with performance nearing that of their commercial counterparts, including Llama-3.2-Vison and Molmo-72B.
+Table 3: Effectiveness of Image-DPO on General VQA benchmarks.
+[…]
+LLaVA-1.5-7B 29.67 37.67 37.7 43.8 12.7 67.3 31.1 49.1 14.8
+LLaVA-1.5-7B + Image-DPO 34.17$\uparrow$ 4.5 39.33$\uparrow$ 1.66 39.79$\uparrow$ 2.09 45.47$\uparrow$ 1.67 14.16$\uparrow$ 1.46 68.45 $\uparrow$ 1.15 32.3$\uparrow$ 1.2 45$\uparrow$ -4.1 12.3$\uparrow$ -2.5
+LLaVA-1.5-13B 35.33 41.5 39.6 44.6 14.8 68.9 36.1 48.3 14.1
+LLaVA-1.5-13B + Image-DPO 38.17$\uparrow$ 2.84 42.5$\uparrow$ 1 42.68$\uparrow$ 3.08 47.37$\uparrow$ 2.77 17.16$\uparrow$ 2.36 70.36$\uparrow$ 1.46 37.5$\uparrow$ 1.4 42.6$\uparrow$ -5.7 11.6$\uparrow$ -2.5
+Cambrian-8B 8.67 32.5 44.6 47.9 19.4 71.5 51.4 14.5 4.7
+Cambrian-8B + Image-DPO 20.83 $\uparrow$ 12.16 39.3 $\uparrow$ 6.83 46.5 $\uparrow$ 1.9 50.2 $\uparrow$ 2.3 20 $\uparrow$ 0.6 72$\uparrow$ 0.5 51.7$\uparrow$ 0.3 11.4 $\uparrow$ -3.1 4.4 $\uparrow$ -0.3
 Do distractor facts really distract?
-In ViLPF setting, we add a distractor fact before the question and explicitly instruct the model to answer with one word.
-Since these facts implicitly suggest incorrect answers for QIAtest, we expected this change to make the questions more suggestive and lower the ViLPFScore, as the distractors would mislead the VLMs.
-But is this true in practice?
-Interestingly, we find that strong models like GPT-4o are not misled by these distractor facts, despite their performance is still significantly lagged behind human accuracy.
-Not only do these models avoid being misled, but surprisingly, they also leverage the distractor factsto improve their answers, leading to a substantial and consistent performance boost when the facts are included.
-We illustrate some examples of this behavior in Figure 4.
-When the distractor fact is provided, GPT-4o performs better on QIAtest compared to when it is not.
-This result is counterintuitive, as the model performs better when its input prompt includes misleading information.
-We hypothesize that the distractor fact may help the model infer the focus of the question, thereby narrowing the search space and ultimately leading to higher accuracy.
-For models with reasonable instruction-following abilities but limited visual reasoning, like LLaVA-1.5-13B (Liu et al., 2023b), the distractor facts behave as expected, often leading the model to incorrect conclusions based solely on textual information.
-This leads to a significant performance drop on ViLPF-Score when distractor facts are present, simultaneously boosting performance on ViLPF-Prior, the language-prior answer.
+In ViLPF setting, we add a distractor fact before the question.
+Since these facts implicitly suggest incorrect answers for QIAtest, we expected this change to make the questions more suggestive and lower the ViLPF-Score, as the distractors would mislead the VLMs.
+Surprisingly, GPT-4o benefits from distractor facts, improving accuracy on QIAtest.
+We hypothesize these facts highlight question focus, narrowing the search space.
+However, weaker models like LLaVA-1.5-13B (Liu et al., 2023b) often get misled by the distractors, hurting their Score but boosting Prior.
 For instance, as shown in Figure 5, with including distractor facts, LLaVA-1.5-13B consistently predicts the distractor fact as the answer.
-However, once the distractors are removed, it can then make the correct prediction.
-In contrast, for models with weaker instruction-following abilities, like Cambrian-8B (Tong et al., 2024), the inclusion of distractor facts significantly increases the difficulty of adhering to explicit instructions, such as providing a single-word answer.
-Specifically, when facts are provided, Cambrian-8B fails to follow the instruction in 62% of the questions, compared to 30% without facts (nearly a 2x increase).
-Upon manual review of these failures, we find that approximately 59% of the responses that deviate from the instruction are still contextually correct, leading to an adjusted accuracy of 47.92%.
-A similar issue is observed with LLaVA-OneVision-72B (Li et al., 2024b), which consistently generates sentences with analysis, when explicitly prompted with “please answer this question with one word".
-This pattern highlights a concerning trend: focusing on improving performance on well-established benchmarks may come at the cost of basic instruction-following abilities, ultimately limiting the practical utility of these models in real-world applications.
-Figure 5: Comparison of benchmark scores under different image transformations. Solid line and dotted line refer to ViLPF-Score and ViLPF-Prior, respectively.
+However, once the distractors are removed, it can then predict correctly.
+For bad instruction-following models like Cambrian-8B (Tong et al., 2024), distractor facts significantly hinder adherence to explicit instructions, such as providing single-word answers.
+With facts, Cambrian-8B fails to follow instructions in 62% of cases, compared to 30% without (a nearly 2x increase).
+Manual review shows 59% of these failures are contextually correct, yielding an adjusted accuracy of 47.92%.
+Similarly, LLaVA-OneVision-72B (Li et al., 2024b) often generates detailed analyses despite explicit single-word prompts.
+This trend highlights a concerning trend: focusing on improving performance on well-established benchmarks may come at the cost of basic instruction-following abilities, ultimately limiting the practical utility of these models in real-world applications.
 How image transformations affect the results?
@@
 Interestingly, GPT-4o, when using degraded images, performs worse in ViLPF-Prior than when no images are used, i.e., GPT-4o (text only) in Table 2.
+Comparison of ViLP with other VQA datasets: To highlight the distinctions between ViLP and existing benchmarks, including Winoground (Thrush et al., 2022), Whoops!
+(Bitton-Guetta et al., 2023a), and HallusionBench (Guan et al., 2023), we conduct a comparative analysis of both their high-level design principles and low-level data formats.
+This comparison incorporates qualitative and quantitative insights, offering a comprehensive understanding of their differences.
+Due to space constraints, detailed analyses are provided in Appendix A.3.
 5.1 Image DPO
-In this section, we examine our proposed Image-DPO (Section 4.1) on both the ViLP  benchmark and general VQA benchmarks.
-Specifically, for self-generating question-image-answer (QIA) pairs, we use COCO (Lin et al., 2014), Text2VQA (Singh et al., 2019b), and Visual Genome (Krishna et al., 2017) as our seed datasets.
-Based on these seed datasets, we instruct the VLM to leverage pretrained image generation and editing models (Liu et al., 2023d; Podell et al., 2023; Brooks et al., 2023b) to generate new images and corresponding QA pairs.
-After generating the QA pairs, we corrupt the images either semantically or at the pixel level, while keeping the QA pairs unchanged, to create the good and bad QIA pairs used for Image DPO training.
-Details of our image generation method are provided in the Appendix C and training details are in the Appendix C.2
+In this section, we evaluate Image-DPO (Section 4.1) on both ViLPand general VQA benchmarks.
+As an ablation baseline, we introduce Text-DPO, which uses the same Question-Image-Answe (QIA) generation process as we used in Image-DPO but applies LLM self-rewarding objective (Yuan et al., 2024) (the standard DPO objective).
+In Text-DPO, good and bad pairs stem from VLM-generated positive and negative answers, while the question and image remain fixed.
+As shown in the right of Figure 2(a), the green box depicts Image-DPO, generating corrupted images via semantic edits, Gaussian blur, and pixelation while keeping the question and answer constant; the purple box illustrates Text-DPO, which fixes the image and varies the answers with associated ratings.
+This setup parallels other VLM self-rewarding work (Zhou et al., 2024a; Deng et al., 2024; Zhou et al., 2024b; Wang et al., 2024c, a).
+For baselines, we compare with VLM self-improvement methods, including SIMA (Wang et al., 2024c), HADPO (Zhao et al., 2023a), and EOS (Yue et al., 2024), by using their publicly available checkpoints.
+Additionally, we train models using the dataset and code provided in RLHF-V (Yu et al., 2023a) and V-DPO (Xie et al., 2024).
+All models use LLaVA-7B for a comprehensive comparisons as many paper only release 7B checkpoints.
+Table 4 shows that Image-DPO achieves the highest across all the metrics.
+Besides, we evaluate the proposed Image-DPO algorithm across three VLM models—Cambrian-8B, LLaVA-1.5-7B, and LLaVA-1.5-13B—using several popular VLM benchmarks that focus on different aspects, including compositionality & biases (NaturalBench (Li et al., 2024a)), general visual reasoning (MM-Vet (Yu et al., 2023c)), and hallucinations (CHAIR (Rohrbach et al., 2018)).
+The results, presented in Table 3, show consistent performance improvements across both datasets and models, further demonstrating the effectiveness of our Image DPO method.
 Table 4: Comparisons of Image-DPO on ViLP.
@@
 LLaVA-1.5-7B 29.67 71.33 37.67 65.67
-LLaVA-1.5-7B + CSR 27.50 64.33 40.84 63.67
-LLaVA-1.5-7B + RLHF-V 29.50 75.00 36.33 65.33
-LLaVA-1.5-7B + Text-DPO 31.34 71.67 37.83 65.67
-LLaVA-1.5-7B + Image-DPO 34.17 77.67 38.67 67.33
-For comparison, we introduce another baseline, Text-DPO, which follows the same question-image-answer generation process as our Image-DPO but applies LLM self-rewarding techniques as outlined in (Yuan et al., 2024).
-In Text-DPO, good and bad pairs are created by sampling positive and negative answers generated by the VLM, while keeping the image and question unchanged.
-The difference between Image-DPO and Text-DPO is illustrated in Figure 6.
-The green box corresponds to Image-DPO, where multiple corrupted images are generated through semantic alterations, Gaussian blurring, and pixelation, while the question and answer remain constant.
-In contrast, Text-DPO (purple-box) maintains the same image and generates multiple answers, each associated with a rating.
-It is worth noting that while Text-DPO shares similarities with other VLM self-rewarding techniques (Zhou et al., 2024a; Deng et al., 2024; Zhou et al., 2024c; Wang et al., 2024b; a), a key distinction is our use of pre-trained image generation models to create diverse images.
-Table 4: Effectiveness of Image-DPO on General VQA benchmarks. The proposed Image-DPO could consistently improve the performance of various VLMs across general VQA benchmarks.
-[…]
-Cambrian-8B 8.67 32.50 44.6 47.9 19.4 14.5 4.7 51.4 11.3
-Cambrian-8B + Image-DPO 11.50 $\uparrow$ 2.83 35.17 $\uparrow$ 2.67 45.68 $\uparrow$ 1.08 49.45 $\uparrow$ 1.55 20.11 $\uparrow$ 0.71 14.3 $\uparrow$ -0.2 4.8 $\downarrow$ 0.1 52.7 $\uparrow$ 1.3 12.0 $\uparrow$ 0.7
-LLaVA-1.5-7B 29.67 37.67 37.7 43.8 12.7 49.1 14.8 31.1 58.6
-LLaVA-1.5-7B + Image-DPO 34.17 $\uparrow$ 4.5 38.67 $\uparrow$ 1 40.68 $\uparrow$ 2.98 46.29 $\uparrow$ 2.49 14.95 $\uparrow$ 2.25 45.4 $\uparrow$ -3.7 12.4 $\uparrow$ -2.4 33.8 $\uparrow$ 2.7 59.65 $\uparrow$ 1.05
-LLaVA-1.5-13B 35.33 41.50 39.6 44.6 14.8 48.30 14.10 36.10 61.9
-LLaVA-1.5-13B + Image-DPO 38.17 $\uparrow$ 2.84 42.5 $\uparrow$ 1 42.68 $\uparrow$ 3.08 47.37 $\uparrow$ 2.77 17.16 $\uparrow$ 2.36 42.6 $\uparrow$ -5.7 11.6 $\uparrow$ -2.5 37.6 $\uparrow$ 1.5 62.32 $\uparrow$ 0.42
-Figure 6: Image-DPO v.s. Text-DPO.
-For comparison, we also report the results of CSR (Zhou et al., 2024b), a VLM self-improvement method that similarly constructs good and bad answers to form training data for DPO.
-Specifically, we use the publicly available checkpoint of the CSR model.
-Additionally, we train a model using RLHF-V (Yu et al., 2023a), a supervised fine-tuning method based on human-annotated data.
-All models are trained on the LLaVA-7B architecture due to computational constraints.
-The results, presented in Table 3, show that our Image-DPO method achieved the highest performance in ViLPF-Score, outperforming Text-DPO, CSR (Zhou et al., 2024b), and the RLHF-V model based on human-annotated data (Yu et al., 2023a).
-On ViLPP-Score, the Image-DPO method achieved the second-best performance.
-Besides, we evaluate the proposed Image-DPO algorithm across three VLM models—Cambrian-8B, LLaVA-1.5-7B, and LLaVA-1.5-13B—using several popular VLM benchmarks that focus on different aspects, including compositionality(NaturalBench (Li et al., 2024a)) hallucinations (CHAIR (Rohrbach et al., 2018)), visual conversation (MM-Vet (Yu et al., 2023c)), and visual reasoning, SEED-Bench (Li et al., 2023a)).
-The results, presented in Table 4, show consistent performance improvements across both datasets and models, further demonstrating the effectiveness of our Image DPO method.
++HADPO (Zhao et al., 2023b) 33.00 74.33 38.50 65.00
++RLHF-V (Yu et al., 2023a) 29.50 75.00 36.33 65.33
++EOS (Yue et al., 2024) 31.33 67.00 38.67 65.67
++SIMA (Wang et al., 2024b) 27.83 68.67 36.17 66.00
++V-DPO (Xie et al., 2024) 29.5 0 72.67 37.83 67.67
++Text-DPO 31.34 71.67 37.83 65.67
++Image-DPO 34.17 75.00 39.33 68.00
+TextDPO on Corrupted Images.
+Does Image-DPO’s improvement stem from the objective itself, or is it merely due to training on more perturbed data?
+To investigate this, we conduct an ablation where we train TextDPO with corrupted images.
+Specifically, for each pair of $(Q,I,A_{w})$ and $(Q,I,A_{l})$ used in Text-DPO, we apply the same corruptions as Image-DPO to $I$ to form $(Q,I^{\prime},A_{w})$ and $(Q,I^{\prime},A_{l})$.
+Results show that LLaVA-v1.5-7B trained with this TextDPO variant achieves a ViLPF-Score of 31, a ViLPP-Score of 37.5, and 30.3 on MMVET.
+In contrast, ImageDPO outperforms it across all metrics, achieving a ViLPF-Score of 34.17, a ViLPP-Score of 39.33, and 32.3 on MMVET.
 6 Conclusion
-In conclusion, we present the ViLPbenchmark to probing the challenge of language bias in Vision-Language Models.
-By utilizing advanced image generation models and designing questions that demand visual cues for accurate responses, our benchmark consists of images that challenge language priors, revealing the limitations of current VLMs, which tend to rely heavily on text.
-Our method, which incorporates self-generated VQA pairs and image corruption for training, has shown promising improvements in enhancing visual reliance, as evidenced by performance gains on models like LLaVA-v1.5 and Cambrian.
+In conclusion, we present the ViLP  benchmark to probe the challenge of visual language bias in Vision-Language Models (VLMs).
+By utilizing advanced image generation models and designing questions that demand visual cues for accurate responses, our benchmark includes images that defy language priors, revealing the limitations of current VLMs.
+Our method, Image-DPO, which incorporates self-generated VQA pairs and image corruption for training, has demonstrated promising improvements in enhancing visual reliance, as evidenced by performance gains on models like LLaVA-v1.5 and Cambrian.
 7 Acknowledgment
@@
 Finally, we greatly appreciate Chris Rockwell for his contributions to our human evaluation studies.
-Appendix
-[…]
-Appendix A More data samples of ViLP
+Appendix A More details and comparisons of our benchmarks
+A.1 More data samples of ViLP
 Figure 6: Randomly sampled data from ViLP.
 Figure 7: Randomly sampled data from ViLP.
+A.2 More details in ViLPbenchmark data generation
+Our proposed dataset introduces Question-Image-Answer (QIA) triplets designed to challenge state-of-the-art Vision-Language Models (VLMs) against visual language priors.
+The construction process combines human-guided and automated efforts to ensure quality and alignment.
+Question-Answer (QA) Generation: Most QA pairs are authored by humans following the design principles in Section 3.1.
+Additionally, candidate QA pairs are generated using models like OpenAI-O1 and Claude-3.5-Sonnet with carefully crafted prompts.
+One example prompts shown in Figure 9.
+These candidates undergo human review, where they are refined or removed to meet our quality standards.
+Image Generation: For each QA pair, we use GPT-4 to generate multiple descriptive image prompts (see Figure 8).
+These prompts are provided to image generation models, such as FLUX and DALL-E 3, to produce candidate images.
+Human reviewers then select the most suitable image or request re-generation as needed to maintain consistency with the QA context.
+Human Review and Testing: At every stage, human reviewers rigorously evaluate the generated outputs to ensure quality, clarity, and challenge level.
+In addition to filtering out low-quality or insufficiently challenging triplets, we dynamically test the QIAs to confirm that they remain intuitive for humans while being difficult for VLMs.
+Cost: The complexity of our data creation process leads to a significant average cost of approximately \$2.50 per QIA triplet in ViLP, excluding human labor costs.
+Task: Using the provided question and possible image-based answers, generate detailed text prompts for image generation.
+Each image prompt should reflect the question’s context and incorporate one of the image-based answers.
+Question: Question
+Image-based Answer:[Answer1, Answer2, Answer3]
+For each possible image-based answer, create an image prompt that describes what the image might look like based on the question.
+Please be creativity.
+For example, if the question asks who is using this mop to clean the floor in the picture?
+and the answer is eraser.
+The image prompt should really describe the image of an eraser uses a mop to clean the floor.
+Format the output strictly as a JSON list, like this example: [ "prompt1": "Image Generation Prompt text here", "prompt2": "Image Generation Prompt text here", "prompt3": "Image Generation Prompt text here", ]
+Figure 8: The prompt we used for generating text-prompt for image generation.
+In the below, I try to propose questions along with three answers where the first answer is corresponding to the question text directly, while the other two are usual and counter-intuitive, which could lead to wrongs of VLMs.
+Please help me generate more Question-3 answer pairs, which are different from what I have provided.
+- All the potential answers should a single world.
+- Help me generate a format where I can direct copy paste into Goole Sheet.
+Also, please a ; between question and each answers.
+- Please be very creative and different from my provided examples - the answer 2 & 3 should be very diverse and different compared to answer 1.
+- Every question contains a statement at the beginning which consists of the answer1 as part of it.
+- Please understand the principles and generate the QA very different from my provided examples Some Examples: - A screwdriver is used for tightening screws.
+From the image, which tool is used to turn screws?
+Screwdriver Hammer Scissors - A pen is a tool used for writing.
+Which object in the image is used to write on paper?
+Pen Hammer Shoe - Clocks are used to measure time.
+Can you identify the item in the image that is used to measure time?
+Clock Spoon Candle - A violin has four strings and is played using a bow.
+According to the image, which musical instrument is being played with a bow?
+Violin Guitar Saxophone - Camels have humps.
+Which animal in the image stores fat in its humps?
+Camel Horse Tiger - Honey is made by bees.
+Which insect in the image produces honey?
+Bee Ant Dragonfly - An anvil is a tool used by blacksmiths.
+What object in the image is used by blacksmiths to forge metal?
+Anvil Fork Wrench - A gavel is used by judges in court.
+From the image, which object symbolizes judicial authority?
+Gavel Hammer Wrench - A syringe is used to inject medicine.
+From the image, which tool is used for administering injections?
+Syringe Scissor Drill - An anchor keeps a ship steady in the water.
+From the image, which item prevents boats from drifting?
+Anchor Spoon Toothbrush - A chainsaw is a power tool for cutting wood.
+What device shown is typically used by lumberjacks to fell trees?
+Chainsaw Blender Stapler
+Figure 9: One prompt we used for potential QAs designs of ViLP
+A.3 Comparisons to other datasets
+Figure 10: Winoground Data Example.
+Our benchmark is different from Winoground, as Winoground focuses on vision-linguistic compositional reasoning.
+Both captions and images are normal and satisfy common linguistic expectations and common sense.
+In this section, we compare our benchmark to other benchmarks, including Winoground (Thrush et al., 2022), Whoops!(Bitton-Guetta et al., 2023a), and HallusionBench(Guan et al., 2023).
+While these datasets are impactful, their evaluation perspectives differ from ours, covering a range from high-level design principles to low-level formats.
+A.3.1 Compare to Winoground (Thrush et al., 2022)
+Figure 11: Example of Whoops! Dataset.Whoops! also has creative images. While unlike ours, its questions are common questions without strong language priors.
+Winoground centers on vision-linguistic compositional reasoning by presenting models with two images and two captions that contain the same words arranged differently.
+The goal is to match each image to its correct caption based on the text’s compositional structures and the visual content (as detailed in the Introduction and Sec.
+3.1 of (Thrush et al., 2022)).
+However, Winoground’s captions do not challenge language priors or introduce out-of-distribution visual information.
+Both captions adhere to common linguistic expectations, and there is no explicit misleading information provided to test resistance to language biases.
+Additionally, most of their images are typical internet images, featuring common visual patterns.
+Figure 12: Example of HallusionBench.
+HallusionBench also has questions which can be answered without images.
+While it is based on facts instead of stereotypes like ours.
+Moreover, its images are limited to Chart, Map, OCR and Table.
+Qualitative Comparison: As shown in Figure 10 consisting of Winoground examples, both captions and images are normal and satisfy common linguistic expectations and common sense.
+The evaluation focuses on whether the model can discern the compositional differences between the two images and two captions, then match them correctly.
+Comparing Figure 10 and Figure 1, 6, 7, You can discern the significant differences among the tested images.
+Quantitative Comparison: Both ViLP; and Winoground benchmarks include paired textual information in their setups.
+In our benchmark, ViLP; Prior QAs and ViLP; Score QAs share the same question but differ in their answers.
+In Winoground, each example has two captions, and the task is to match each caption to its correct image.
+Setting.
+To demonstrate the differences, we use GPT to evaluate the commonness of these paired textual components.
+Specifically, GPT-4o rates the oddity of scenarios described in texts on a scale from 1 (very rare) to 10 (very common).
+The resulting scores are then compared.
+Results.
+In our benchmark, Prior QAs scored 9.37, indicating that these answers are designed to align with language priors and are highly common.
+Score QAs scored 1.65, showing that these QA pairs are rare, making them difficult to infer without the corresponding visual information.
+Notably, Prior and Score QAs share the same question but differ in their answers, and this significant contrast in scores showcases how we inject strong language priors to test a model’s vulnerability to linguistic distractions.
+By comparison, Winoground’s two captions scored 8.05 and 8.08, indicating two primary observations: (1) both captions align well with language priors, which means Winoground does not challenge language priors or evaluate out-of-distribution scenarios; (2) the minimal score difference between the two captions confirms there is no significant variance in language priors, as examining how VLM models react to different language priors is beyond the scope of Winoground.
+In contrast, that aspect is precisely our focus.
+A.3.2 Compare to Whoops! (Bitton-Guetta et al., 2023a)
+Whoops!
+is designed to evaluate a model’s ability to detect weirdness in images, emphasizing tasks where images depict unusual or nonsensical scenarios.
+It heavily relies on common sense reasoning, requiring models to recognize visual elements and then identify subtle inconsistencies among them.
+For example, for the lit candle inside a tightly sealed glass jar on the homepage, models must realize that “a candle needs a constant oxygen supply to burn, which would not exist in a sealed jar", making a burning candle inside a sealed jar unlikely.
+This benchmark thus focuses on common-sense reasoning rather than challenging visual language priors.
+Qualitative Comparison
+Although Whoops!
+also includes creative, out-of-distribution images, it does not focus on using language priors to test a model’s susceptibility to linguistic distraction, as our benchmark does.
+In its QA mode (comparable to our task), the questions are straightforward and lack strong language priors.
+Some examples can be found in Figure 11.
+Additionally, Whoops!
+uses open-ended questions, offering greater freedom in answers while introducing potential ambiguity or divergence in responses.
+Quantitative Comparison
+Setting.
+Unlike Winoground and our benchmark, Whoops!
+does not provide control groups or textual components for comparison.
+To measure language priors, we analyze the suggestiveness of questions by evaluating GPT-4o’s certainty when answering them without any visual context.
+A more suggestive question typically yields more determined and confident responses, whereas a less suggestive question produces more varied answers.
+We calculate how many unique answers GPT-4o provides over five attempts at temperature 1.0 to promote randomness.
+Semantic differences are normalized to exclude synonyms.
+Results.
+We find that Whoops!
+questions produce an average of 2.58 unique answers (out of five attempts) with a standard deviation of 1.48.
+For our benchmark, without facts, GPT-4o provides an average of 1.53 unique answers (std 0.94), and with facts, 1.10 unique answers (std 0.42).
+Although both benchmarks use creative images, these results indicate that Whoops!
+questions remain more general and do not push GPT-4o toward stereotypical responses.
+In contrast, our benchmark deliberately uses suggestive questions to elicit stereotype-consistent answers, reflecting our emphasis on testing language priors.
+A.3.3 Compare to HallusionBench (Guan et al., 2023)
+HallusionBench has two components: Visual Dependent, which focuses on testing models’ general visual reasoning skills, and Visual Supplement, which examines a model’s visual reasoning in tandem with its parametric memory.
+The Visual Supplement part is related to our benchmark, as its questions, like ours, can be answered without visual information.
+However, the key difference lies in their design.
+HallusionBench questions rely on parametric memory and strict factual knowledge (e.g., “Which country has the most gold medals in the 2024 Olympics?”), whereas our benchmark questions are based on common stereotypes (e.g., “A soccer ball is round.”).
+This distinction significantly constrains the diversity of HallusionBench’s questions—its official release contains only 50 question pairs—while our approach allows for a broader range of scenarios, including texture, shape, conceptual combinations, hallucinated elements, and proverb-based contexts.
+Moreover, HallusionBench focuses on a narrow set of image types (charts, tables, maps, and OCR).
+It tests reasoning through subtle modifications (like changing one digit in a chart), which is fundamentally different from our approach of creating diverse visual contexts.
+Figure 12 showcases typical HallusionBench examples, illustrating the limited scope of both images and QAs, whereas our benchmark includes a variety of out-of-distribution images and more suggestive QAs (see Figure 1, 6, 7).
+A.4 Human Study
+For human evaluation, we hired Ph.D.-level candidates to participate in testing.
+They were asked to answer questions with a single image provided each time, and the QIAs were randomly shuffled to avoid any sequential context.
+To ensure efficiency, we conducted an oral test instead of a written one, recording their responses.
+After the test, we updated the synonym sets for the QIAs based on their answers.
 Appendix B Image-DPO Mathematical Details
@@
 This stage produces a large number of new question-image-answer (QIA) triplets.
-In the second stage, we apply three types of image corruptions—Gaussian blurring, pixelization, and semantic editing—to generate good-bad QIA pairs, denoted as $I_{w}$ (good) and $I_{l}$ (bad).
+In the second stage, we apply three types of image corruptions—Gaussian blurring, pixelation, and semantic editing—to generate good-bad QIA pairs, denoted as $I_{w}$ (good) and $I_{l}$ (bad).
 Details of the hyperparameters used in the experiments are provided at the end of this section.
@@
 The used text prompt is shown in Figure 16.
-Subsequently, pre-trained models such as Stable Diffusion XL Podell et al.
-(2023), Instruct-Pix2Pix Brooks et al.
-(2023b), and Grounded-SAM Ren et al.
-(2024) are employed to either modify existing images or generate entirely new ones.
+Subsequently, pre-trained models such as Stable Diffusion XL (Podell et al., 2023), Instruct-Pix2Pix (Brooks et al., 2023b), and Grounded-SAM (Ren et al., 2024) are employed to either modify existing images or generate entirely new ones.
 The altered or newly created images, along with the instructions that guided their generation, are then used by the same VLMs to produce corresponding question-answer pairs (QAs) based on the text prompt shown in Figure 17.
@@
 To facilitate this, we use an additional text prompt (Figure 18) after the VLM generates the initial instructions (the pink region of Figure 15).
-To provide a better understanding of our generated QIAs, we randomly sampled and listed some examples of the generated QIA data, as shown in Figures 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, and 29.
+To provide a better understanding of our generated QIAs, we randomly sampled and listed some examples of the generated QIA data, as shown in Figures 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, and 33.
 Figure 15: Randomly sampled generation examples. Our proposed data generation pipeline utilizes pretrained model to generate diverse new images from existing seed image datasets.
@@
 Your recommendations should help in understanding the potential of the image and exploring creative possibilities.
-Expected Response Format: Item Number: 1 Tool Used: [Specify the tool - Stable Diffusion or InstructPix2Pix or GroundingDINO] Text Prompt for Processing: [Detailed description of the task or transformation to be performed.
-For image generation, please provide complete description based on the understanding of the provided images, since we only feed text prompt for this task.] Item Number: 2 Tool Used: [Specify the tool - Stable Diffusion or InstructPix2Pix or GroundingDINO] Text Prompt for Processing: [Detailed description of the task or transformation to be performed.
-For image generation, please provide complete description based on the understanding of the provided images, since we only feed text prompt for this task.] Item Number: 3 Tool Used: [Specify the tool - Stable Diffusion or InstructPix2Pix or GroundingDINO] Text Prompt for Processing: [Detailed description of the task or transformation to be performed.
+Expected Response Format: Item Number: 1
+Tool Used: [Specify the tool - Stable Diffusion or InstructPix2Pix or GroundingDINO]
+Text Prompt for Processing: [Detailed description of the task or transformation to be performed.
+For image generation, please provide complete description based on the understanding of the provided images, since we only feed text prompt for this task.]
+Item Number: 2
+Tool Used: [Specify the tool - Stable Diffusion or InstructPix2Pix or GroundingDINO]
+Text Prompt for Processing: [Detailed description of the task or transformation to be performed.
+For image generation, please provide complete description based on the understanding of the provided images, since we only feed text prompt for this task.]
+Item Number: 3
+Tool Used: [Specify the tool - Stable Diffusion or InstructPix2Pix or GroundingDINO]
+Text Prompt for Processing: [Detailed description of the task or transformation to be performed.
 For image generation, please provide complete description based on the understanding of the provided images, since we only feed text prompt for this task.]
@@
 Each question should be followed by a comprehensive answer, providing depth and context.
-Expected Multiple Response Format: Item Number: 1 Question: [Propose a unique and insightful question based on the descriptions and the images.] Answer: [Provide a comprehensive answer to the proposed question.] Item Number: 2 Question: [Propose a unique and insightful question based on the descriptions and the images.] Answer: [Provide a comprehensive answer to the proposed question.] Please ensure each question-answer pair is well-defined and informative.
-Please provide at least 5 question-answer pairs based on the input provided.
+Expected Multiple Response Format:
+Item Number: 1
+Question: [Propose a unique and insightful question based on the descriptions and the images.]
+Answer: [Provide a comprehensive answer to the proposed question.]
+Item Number: 2
+Question: [Propose a unique and insightful question based on the descriptions and the images.]
+Answer: [Provide a comprehensive answer to the proposed question.]
+Please ensure each question-answer pair is well-defined and informative. Please provide at least 5 question-answer pairs based on the input provided.
 Figure 17: The prompt for single-image QAs. We ask the VLM itself to generate single-image QAs based on the generated images by pre-trained models.
 Analyze the provided image and its accompanying modification instruction to identify the removed object description, the new object description, and the new image description.
-Modification Instructions: <Text Prompt for Processing> Expected Multiple Response Format: Item Number: 1 Removed Object Description: [Brief description of the object to be detected and removed] New Object Description: [Description of a new, different object to replace the removed one] New Image Description: [Description of the image after each object’s removal, focusing on changes and remaining elements] Item Number: 2 Removed Object Description: [Brief description of the object to be detected and removed] New Object Description: [Description of a new, different object to replace the removed one] New Image Description: [Description of the image after each object’s removal, focusing on changes and remaining elements]
+Modification Instructions: <Text Prompt for Processing> Expected Multiple Response Format:
+Item Number: 1
+Removed Object Description: [Brief description of the object to be detected and removed]
+New Object Description: [Description of a new, different object to replace the removed one]
+New Image Description: [Description of the image after each object’s removal, focusing on changes and remaining elements]
+Item Number: 2
+Removed Object Description: [Brief description of the object to be detected and removed]
+New Object Description: [Description of a new, different object to replace the removed one]
+New Image Description: [Description of the image after each object’s removal, focusing on changes and remaining elements]
 Figure 18: The prompt for instruction generation of Grounded-SAM. We ask the VLM to generate designated instructions to use Grounded-SAM.
@@
 In brief, the data generation pipeline outlined in Appendix C.1 utilizes VLMs in conjunction with pre-trained image models to generate or modify images and create corresponding question-answer pairs.
-This process results in a collection of QIA triplets, as illustrated in the Figures 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, and 29.
-After generating the QIA triplets, we apply three image corruption methods—Gaussian blurring, pixelization, and semantic editing—to create good-bad QIA pairs for ImageDPO training (Section 4.1), while keeping the QA components unchanged.
+This process results in a collection of QIA triplets, as illustrated in the Figures 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, and 33.
+After generating the QIA triplets, we apply three image corruption methods—Gaussian blurring, pixelation, and semantic editing—to create good-bad QIA pairs for ImageDPO training (Section 4.1), while keeping the QA components unchanged.
 For Gaussian blur, we use a kernel size of 40 for Cambrian and 80 for LLaVA, as the larger kernel size showed better performance for LLaVA.
-For pixelization, we apply block sizes of 32 for Cambrian and 64 for LLaVA.
+For pixelation, we apply block sizes of 32 for Cambrian and 64 for LLaVA.
 For these two types of corruption, we utilize data generated by all three pre-trained models: Stable Diffusion, Instruct-Pix2Pix, and Grounded-SAM.
@@
 Figure 33: Randomly sampled single-image QA generated by our data generation pipeline.
-Appendix D More details and comparisons of our benchmarks
-D.1 Our benchmark data generation
-Our proposed dataset introduces QIA triplets meticulously designed to challenge state-of-the-art Vision-Language Models (VLMs) by mitigating language priors.
-The construction process combines human-guided and automated efforts to ensure quality and alignment across all components.
-Question-Answer (QA) Generation: Questions and answers are partially authored by humans, adhering to the design principles outlined in Section 3.1.
-Additionally, candidate QA pairs are generated using models like OpenAI-O1 and Claude-3.5-Sonnet with carefully crafted prompts.
-One example prompts shown in Figure 30.
-These candidates undergo rigorous human review, where they are refined or filtered to meet our quality standards.
-Image Generation: To align with the QA pairs, descriptive and diverse image prompts are first generated using GPT-4 (see Figure 31).
-These prompts are then input into advanced image generation models, such as FLUX and DALL-E 3, to produce candidate images.
-Multiple images are generated per QA pair, and human reviewers carefully select the most suitable image or request re-generation to ensure alignment with the QA context.
-Human Review and Testing: At every stage, human reviewers rigorously evaluate the generated outputs to ensure quality, clarity, and challenge level.
-In addition to filtering out low-quality or insufficiently challenging triplets, we also test the QIAs with users to confirm that they remain intuitive for humans while being difficult for VLMs.
-Overall, the complexity of creating our dataset results in a high average cost of generating a single QIA triplet in ViLP, amounting to approximately \$7.50, excluding human labor costs.
-In the below, I try to propose questions along with three answers where the first answer is corresponding to the question text directly, while the other two are usual and counter-intuitive, which could lead to wrongs of VLMs.
-Please help me generate more Question-3 answer pairs, which are different from what I have provided.
-- All the potential answers should a single world.
-- Help me generate a format where I can direct copy paste into Goole Sheet.
-Also, please a ; between question and each answers.
-- Please be very creative and different from my provided examples - the answer 2 & 3 should be very diverse and different compared to answer 1.
-- Every question contains a statement at the beginning which consists of the answer1 as part of it.
-- Please understand the principles and generate the QA very different from my provided examples Some Examples: - A screwdriver is used for tightening screws.
-From the image, which tool is used to turn screws?
-Screwdriver Hammer Scissors - A pen is a tool used for writing.
-Which object in the image is used to write on paper?
-Pen Hammer Shoe - Clocks are used to measure time.
-Can you identify the item in the image that is used to measure time?
-Clock Spoon Candle - A violin has four strings and is played using a bow.
-According to the image, which musical instrument is being played with a bow?
-Violin Guitar Saxophone - Camels have humps.
-Which animal in the image stores fat in its humps?
-Camel Horse Tiger - Honey is made by bees.
-Which insect in the image produces honey?
-Bee Ant Dragonfly - An anvil is a tool used by blacksmiths.
-What object in the image is used by blacksmiths to forge metal?
-Anvil Fork Wrench - A gavel is used by judges in court.
-From the image, which object symbolizes judicial authority?
-Gavel Hammer Wrench - A syringe is used to inject medicine.
-From the image, which tool is used for administering injections?
-Syringe Scissor Drill - An anchor keeps a ship steady in the water.
-From the image, which item prevents boats from drifting?
-Anchor Spoon Toothbrush - A chainsaw is a power tool for cutting wood.
-What device shown is typically used by lumberjacks to fell trees?
-Chainsaw Blender Stapler
-Figure 30: One prompt we used for potential QAs designs of ViLP
-Task: Using the provided question and possible image-based answers, generate detailed text prompts for image generation.
-Each image prompt should reflect the question’s context and incorporate one of the image-based answers.
-Question: Question Image-based Answer:[Answer1, Answer2, Answer3] For each possible image-based answer, create an image prompt that describes what the image might look like based on the question.
-Please be creativity.
-For example, if the question asks who is using this mop to clean the floor in the picture?
-and the answer is eraser.
-The image prompt should really describe the image of an eraser uses a mop to clean the floor.
-Format the output strictly as a JSON list, like this example: [ "prompt1": "Image Generation Prompt text here", "prompt2": "Image Generation Prompt text here", "prompt3": "Image Generation Prompt text here", ]
-Figure 31: The prompt we used for generating text-prompt for image generation.
-D.2 Comparisons to other datasets
-Here we compare our benchmark to other popular benchmarks.
-All these compared datasets are well-designed and impactful.
-However, they have different evaluation focus compared to ours, from high-level design principles to low-level formats.
-D.2.1 Compare to Winoground Thrush et al. (2022)
-Winoground focuses on vision-linguistic compositional reasoning by presenting models with two images and two captions containing identical sets of words arranged differently.
-The goal is to match images to captions based on compositional structures in the text and visual content (as detailed in the Introduction and Sec 3.1 of the referenced paper).
-However, the captions in Winoground do not challenge language priors or introduce out-of-distribution information.
-Both captions remain consistent with common linguistic expectations, and no explicit misleading information is provided to test resistance to language biases.
-Figure 32: Winoground Data Example.
-Our benchmark is different from Winoground, as Winoground focuses on vision-linguistic compositional reasoning.
-Both captions and images are normal and satisfy common linguistic expectations and common sense.
-Qualitative Comparison As shown in Figure 32 consisting of Winoground examples, both captions and images are normal and satisfy common linguistic expectations and common sense.
-The evaluation focus is on whether or not the model can tell the compositional difference between the two images and two captions, and match them correctly.
-Quantitative Comparison Both ViLP  and Winoground benchmarks include paired textual information in their settings.
-In our benchmark, we have ViLP  Prior QAs and ViLP  Score QAs, which share the same question but differ in their answers.
-In Winoground, for each example, there are two captions, and the task is to match the correct image to the correct caption.
-To demonstrate the differences, we use GPT to evaluate the commonness of these paired textual components.
-Specifically, GPT-4o rates the oddity of scenarios described in texts on a scale from 1 (very rare) to 10 (very common), and these scores are compared.
-In our benchmark, Prior QAs scored 9.37, indicating answers designed to align with language priors are easily guessed and highly common.
-Score QAs scored 1.65, showing that these QA pairs are significantly rare, which are difficult to infer without the corresponding visual information.
-Notably, Prior and Score QAs share the same question but differ in their answers.
-The stark difference in scores demonstrates our injection of strong language priors, used to test the model’s susceptibility to linguistic distractions.
-For comparison, Winoground’s two captions yielded scores of 8.05 and 8.08, revealing two major differences:
-Both captions align perfectly with language priors (at least according to GPT-4o), indicating that Winoground does not challenge language priors or evaluate out-of-distribution scenarios The minimal score difference between the two captions indicates no significant variance in language priors, as exploring how VLM models behave under varying language priors is not within the scope of this benchmark.
-In contrast, this is precisely the focus of our benchmark.
-D.2.2 Compare to Whoops! Bitton-Guetta et al. (2023a)
-Whoops!
-is designed to evaluate a model’s ability to detect weirdness in images, focusing on tasks where images depict unusual or nonsensical scenarios.
-It highly emphasizes common sense reasoning: models are expected to recognize visual elements and then reason the subtle contrast between these elements to identify the oddity.
-For example, for A lit candle is sitting inside a tightly sealed glass jar example in the homepage, models are expected to realize that A candle needs a constant supply of oxygen to burn, which does not exist in a sealed bottle, so it is unlikely to see a burning candle inside a sealed bottle.
-This benchmark emphasizes common-sense reasoning instead of defying language priors.
-Figure 33: Example of Whoops! Dataset.Whoops! also has creative images. While unlike ours, its questions are common questions without strong language priors.
-Qualitative Comparison
-Although its images are creative and out of distribution like ours, this benchmark does not focus on using language priors to test the model’s susceptibility to linguistic distractions as our benchmark does.
-Especially, in their QA mode(which is the same task as our benchmark), the questions are normal questions without strong language priors like ours.
-Some examples can be found in Figure 33.
-Also, it uses a format of open-ended questions.
-This format allows higher freedom in QA, while adding ambiguity and even offs in the answers.
-Quantitative Comparison
-Unlike Winoground and our benchmark, Whoops!
-does not provide control groups or textual components for comparison.
-To measure language priors, we analyze the suggestiveness of questions by evaluating how determined and confident GPT-4o answers them (without visual context).
-A more suggestive question is expected to yield more determined and confident answers, while a less suggestive question is expected to yield more diverse answers.
-For this, we calculate the number of unique answers given by GPT-4o over five attempts while setting its temperature to 1.0 for higher randomness.
-Semantic differences were normalized to exclude synonyms.
-Interestingly, we find questions from Whoops!
-yield 2.58 unique answers (in 5 attempts) on average with std 1.48.
-For our benchmark, without facts, GPT-4o gives 1.53 unique answers (in 5 attempts) on average with std 0.94.
-With facts, GPT-4o gives 1.10 unique answers (in 5 attempts) on average with std 0.42.
-While both benchmarks include creative images, these results show that Whoops!
-questions adhere to conventional styles, which do not prompt GPT-4o to stereotype its answers.
-In contrast, our benchmark’s suggestive questions are intentionally designed to elicit stereotype-averse responses, aligning with our design principles.
-D.2.3 Compare to HallusionBench Guan et al. (2023)
-HallusionBench consists of 2 parts: Visual Dependent (which focuses on testing a model’s general visual reasoning skills) and Visual Supplement (which evaluates a model’s visual reasoning ability and the balance between parametric memory and image context).
-The Visual Supplement part is related to our benchmark, as its questions, like ours, can be answered without visual information.
-However, the key difference lies in their design.
-HallusionBench questions rely on parametric memory and strict factual knowledge, e.g., Which country has the most gold medals in the 2024 Olympics?, whereas our questions are based on stereotypes, e.g., A soccer ball is round.
-This design in HallusionBench significantly constrains the scope of questions and the benchmark’s size, as the official release includes only 50 question pairs.
-Additionally, it is unclear whether VLM models are expected to rely on their factual knowledge or the inputted, modified chart to answer these questions, as the altered images contradict real-world facts.
-Furthermore, all of HallusionBench’s images fall into specific categories: chart, table, map, and OCR, representing a very narrow range of visual information.
-While the benchmark claims to test a model’s reasoning ability by making subtle changes to these images (e.g., altering a single digit in a chart), reading and understanding such images is already challenging, even without modifications.
-And these minor changes further complicate the task, and makes the results hard to interpret.
-Figure 34: Example of HallusionBench.
-HallusionBench also has questions which can be answered without images.
-While it is based on facts instead of stereotypes like ours.
-Moreover, its images are limited to Chart, Map, OCR and Table.
-We show representative images of HallusionBench in this image, which clearly illustrates the very narrow range of image and QA types compared to our benchmark.
-The modification of the image (from green boxed image to red one) is hard to recognize even for humans.
-D.3 Human Study
-For human evaluation, we hired Ph.D.-level candidates to participate in testing.
-They were asked to answer questions with a single image provided each time, and the QIAs were randomly shuffled to avoid any sequential context.
-To ensure efficiency, we conducted an oral test instead of a written one, recording their responses.
-After the test, we updated the synonym sets for the QIAs based on their answers.
-Every test session lasted approximately 1.5 hours, and all the participants expressed confidence in their answers.

### v4

@@
 Moreover, we theoretically show that our objective optimizes an upper bound of the RLHF objective (Ouyang et al., 2022).
-The proposed ViLP dataset, Image-DPO code, and our synthetic data appear in https://huggingface.co/ViLP.
+The proposed ViLP dataset, Image-DPO code, and our synthetic data appear in https://vilp-team.github.io/.
 Figure 1: Sample data from ViLP.
@@
 Please refer to Appendix A.1 for more data samples from ViLP.
+We investigate the impact of image styles in Appendix A.3, where we generate more realistic images using 4o image generation.
 Furthermore, we include both qualitative and quantitative comparison results with Winoground (Thrush et al., 2022), Whoops!(Bitton-Guetta et al., 2023a), and HallusionBench(Guan et al., 2023) in Appendix A.4.
@@
 Comprehensive comparisons with existing datasets appear in Appendix A.4.
-Vision Language Models and Language Priors: Multimodal reasoning is crucial for machine intelligence, with VLMs integrating visual perception, text reasoning, instruction following, and generation for complex tasks (Tan & Bansal, 2019; Li et al., 2019; Kim et al., 2021; Wang et al., 2021b, a; Alayrac et al., 2022; Li et al., 2023b; Chen et al., 2022; Jia et al., 2021; Shen et al., 2021; Singh et al., 2021; Liu et al., 2023c, a; Zhao et al., 2023a; Chen et al., 2023; Zhu et al., 2024; Li et al., 2024c; Dai et al., 2023; Li et al., 2024c; Yu et al., 2024; Dai et al., 2024; Deitke et al., 2024).
+Vision Language Models and Language Priors: Multimodal reasoning is crucial for machine intelligence, with VLMs integrating visual perception, text reasoning, instruction following, and generation for complex tasks (Tan & Bansal, 2019; Li et al., 2019; Kim et al., 2021; Wang et al., 2021b, a; Alayrac et al., 2022; Li et al., 2023b; Chen et al., 2022; Jia et al., 2021; Shen et al., 2021; Singh et al., 2021; Liu et al., 2023c, a; Zhao et al., 2023a; Chen et al., 2023; Zhu et al., 2024b; Li et al., 2024c; Dai et al., 2023; Li et al., 2024c; Yu et al., 2024; Dai et al., 2024; Deitke et al., 2024).
 Inspired by the success of large language models (Brown et al., 2020; OpenAI, 2023a; Touvron et al., 2023a, b; Chiang et al., 2023) and pre-trained visual encoders (Radford et al., 2021; Desai & Johnson, 2020; Caron et al., 2021; Chen et al., 2024), many recent methods leverage relatively small vision-language paired datasets (Liu et al., 2024a; Tong et al., 2024) to fine-tune connectors between LLMs and visual backbones (Liu et al., 2024a).
@@
 This approach extends to VLMs by generating new answers for DPO training (Zhou et al., 2024a; Deng et al., 2024; Zhou et al., 2024b; Wang et al., 2024c, a; Yue et al., 2024; Liu et al., 2024b; Xiao et al., 2024).
-Our work aligns with these self-rewarding VLMs but differs in two key ways: (1) our proposed Image-DPO generates multiple images for a single question-image pair (rather than multiple answers); (2) rather than relying solely on existing images, Image-DPO pipeline creates diverse new images using pre-trained generative models (SDXL (Podell et al., 2023), GroundedSAM (Ren et al., 2024), InstructPix2Pix (Brooks et al., 2023a)).
+Our work aligns with these self-rewarding VLMs but differs in two key ways: (1) our proposed Image-DPO generates multiple images for a single question-image pair (rather than multiple answers); (2) rather than relying solely on existing images (Zhu et al., 2024a), Image-DPO creates diverse new images using pre-trained generative models (SDXL (Podell et al., 2023), GroundedSAM (Ren et al., 2024), InstructPix2Pix (Brooks et al., 2023a)).
 Furthermore, Image-DPO deliberately corrupts images to produce multiple degraded versions that serve as rejected data in DPO training.
-Concurrent works (Wang et al., 2024a; Xie et al., 2024) explore similar methodologies but lack a benchmark to verify enhanced visual focus, fail to establish theoretical connections between their proposed objective and DPO, and utilize limited image transformations (e.g., only randomly cropping).
+Concurrent works (Wang et al., 2024a; Xie et al., 2024) explore similar methods but lack a benchmark to verify enhanced visual focus, fail to establish theoretical connections between their proposed objective and DPO, and utilize limited image transformations (e.g., only randomly cropping).
 In contrast, we introduce ViLP (Section 3) to assess visual reasoning and provide theoretical foundations (Appendix B), alongside multi-category image corruptions (semantic editing, Gaussian blurring, pixelation).
@@
 First, text-only inference ensures that each question can be answered with high confidence using textual clues alone.
-Second, Visual Influence requires that the correct answer—sometimes contradicting common sense—only emerges once an out-of-distribution image is considered.
+Second, visual inference requires that the correct answer—sometimes contradicting common sense—only emerges once an out-of-distribution image is considered.
 By forcing models to integrate both textual and visual information, ViLP reveals whether they truly engage in visual reasoning or merely rely on memorized patterns.
-Mathematically, let $Q$ be a question, $I$ an image, and $A={a_{\text{prior}},\dots,a_{\text{test}},\dots}$ the set of possible answers.
+Mathematically, let $Q$ be a question, $I$ an image, and $A=\{a_{\text{prior}},\dots,a_{\text{test}},\dots\}$ the set of possible answers.
 We define $P(a\mid Q)$ as the probability of answer $a$ given $Q$ alone and $P(a\mid Q,I)$ as the probability given both $Q$ and $I$.
@@
 $P(a_{\text{test}}\mid Q)\leq\delta_{4},\,P(a_{\text{prior}}\mid Q,I)\leq\delta_{4}$ (3)
-In constructing ViLP, we use human cognition prior $P_{\text{human}}$ as guidance for its design.
-We evaluated the behavior of the learned priors of VLMs/LLMs by benchmarking $P_{\theta}$ against the human prior $P_{\text{human}}$.
+In designing ViLP, we leverage the human cognition prior $P_{\text{human}}$ as our guiding principle, ensuring each QIA configuration aligns with typical human expectations while requiring visual evidence to override strong textual assumptions.
+We then compare the learned priors of VLMs and LLMs, denoted $P_{\theta}$, $P_{\text{human}}$ to evaluate whether these models genuinely engage in visual reasoning rather than relying on memorized patterns.
 3.2 Question-Image-Answer Generation
@@
 We incorporate substantial human input and leverage advanced LLMs such as OpenAI-o1 and Claude-3.5-Sonnet to ensure alignment with all the criteria.
-More details are provided in Appendix A.2.
+More details, including text prompts and average cost, are provided in Appendix A.2.
+Note that as more advanced image generative models—such as the recently introduced 4o image generation—become available, we anticipate generating increasingly abundant and high-quality data for our benchmark, yet future updates will remain consistent with the dataset construction criteria outlined in Section 3.1.
 (a) Image-DPO vs.
@@
 Rigorous human review ensures that all final QIA triplets are clear and interpretable, as reflected by our human evaluation results in Table 2.
+Besides, in Appendix A.3, we investigate the impact of image styles by generating more realistic images via 4o image generation and comparing them to those produced by DALL·E-3 (Ramesh et al., 2021) and Flux.
+We find that realistic images can increase the difficulty of the task, highlighting their importance for future studies.
 3.3 Dataset Evaluation
-All of our questions are designed to be answered with a single word, as this design greatly simplifies evaluation protocol compared to assessing full sentences, without requiring LLMs for evaluation.
+All of our questions are designed to elicit single-word answers, an approach that is more efficient and more reliable than sentence-based evaluations that rely on LLM judgment.
+By avoiding using LLM, we reduce API fees, computational overhead, and the risk of occasional inaccuracies due to incorrect model reasoning.
 We explicitly instruct the model to provide a single-word answer, and we evaluate the correctness of each response using a binary system.
-To ensure a fair evaluation, we devote significant effort to building a comprehensive set of synonyms and plural for each answer to detect other valid alternative answers.
+To ensure a fair evaluation, we devote significant efforts to building a comprehensive set of synonyms and plural for each answer to detect other valid alternative answers.
 This ensures that the model is only penalized for actual errors, not for providing synonymous or alternative correct responses.
@@
 4.1 Objective
-Figure 3: Qualitative examples. We show the results from GPT-4o, Claude-3.5-Sonnet, Gemini-1.5-Pro, and Llama-3.2-Vision-90B for some challenging cases.
+Figure 3: Qualitative examples.
+We show the results from GPT-4o, Claude-3.5-Sonnet, Gemini-1.5-Pro, and Llama-3.2-Vision-90B for some challenging cases.
+Please refer to Appendix A.6 for more failure case analysis.
 Existing approaches for VLM self-improvement follow the way used in DPO paper (Rafailov et al., 2024), where the model is trained to distinguish between good and bad answers for a fixed image and question (Figure 2(a) right).
@@
 The triplet $(Q,I_{l},A)$ forms a degraded question-image-answer pair compared to $(Q,I_{w},A)$.
-We train the model to distinguish between good and bad triplets using the objective below, where $\theta$ is the target VLM, $\theta_{\text{ref}}$ is the reference VLM (typically an earlier version of $\theta$), $S$ is the dataset of good and bad triplets, $\sigma$ is the sigmoid function, and $\alpha$ is a scaling factor.
-$L(\theta,\theta_{\text{ref}})$ $=-\mathbb{E}_{Q,I_{w},I_{l},A\sim S}\Bigl{[}$ (4)
-$\quad\log\sigma\Bigl{(}\alpha\frac{\pi_{\theta}(A\mid Q,I_{w})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{w})}\;-\;\alpha\frac{\pi_{\theta}(A\mid Q,I_{l})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{l})}\Bigr{)}\Bigr{]}$
+We train the model to distinguish between good and bad triplets using the objective 4, where $\pi_{\theta}$ is the target VLM, $\pi_{\text{ref}}$ is the reference VLM (typically an earlier version of $\pi_{\theta}$), $S$ is the dataset of good and bad triplets, $\sigma$ is the sigmoid function, and $\alpha$ is a scaling factor.
+$L(\pi_{\theta},\pi_{\text{ref}})$ $=-\mathbb{E}_{Q,I_{w},I_{l},A\sim S}\Bigl{[}$ (4)
+$\quad\log\sigma\Bigl{(}\alpha\frac{\pi_{\theta}(A\mid Q,I_{w})}{\pi_{{\text{ref}}}(A\mid Q,I_{w})}\;-\;\alpha\frac{\pi_{\theta}(A\mid Q,I_{l})}{\pi_{{\text{ref}}}(A\mid Q,I_{l})}\Bigr{)}\Bigr{]}$
 Intuitively, since the textual inputs and outputs are identical in both good and bad cases, the gradients of this objective push the model to rely more on the vision branch, driving a shift in gradient direction when processing normal images $I_{w}$ compared to corrupted images $I_{l}$ (Figure 15).
@@
 LLaVA-OneVision-72B 1.67 3.00 5.22 11.67
-We introduce ViLP,  a benchmark of 300 questions, each with three unique images and corresponding answers (QIAprior and two QIAtest), yielding 900 QIAs.
-QIAprior aligns with common language priors (easily inferred from text alone), while QIAtest challenges these priors, requiring visual reasoning.
-ViLP features two evaluation settings: ViLPF (with distractor facts) and ViLPP (pure questions).
-We report average accuracy on QIAtest (Score) and on QIAprior (Prior) in Table 2.
+We introduce ViLP, a new benchmark comprising 300 questions.
+Each question is paired with three unique images and their corresponding answers—one QIAprior and two QIAtest—for a total of 900 QIAs.
+The QIAprior examples (300 in total) align with common language priors (i.e., they can usually be answered correctly by relying on textual cues alone).
+In contrast, the QIAtest examples (600 in total) challenge these priors by requiring visual reasoning.
+ViLP features two evaluation settings:
+  • ViLPF, where both distractor facts and the questions are provided;
+  • ViLPP, where only the questions themselves are given (i.e., no distractor facts).
+We report two metrics in Table 2: average accuracy on QIAtest (noted as Score) and average accuracy on QIAprior (noted as Prior). Our benchmark emphasizes the performance in Score.
 Figure 4: Comparison of benchmark scores under different image transformations. Solid line and dotted line refer to ViLPF-Score and ViLPF-Prior, respectively.
@@
 Participants achieved nearly perfect accuracy on ViLPF-Prior and over 98% on ViLPF-Score and ViLPP-Score, confirming that our question-image-answer combinations are unambiguous for human interpretation.
-Notably, despite QIAtest being designed as out-of-distribution examples, humans were still able to correctly distinguish them
+Notably, despite QIAtest being designed as out-of-distribution examples, humans were still able to correctly distinguish them.
 Humans performed slightly better on ViLPF-Prior when distractor facts were provided, as they could easily identify that these facts aligned with the correct answers.
@@
 Notably, it is encouraging to see that some open-source models achieved over 60% accuracy on ViLPF-Score, with performance nearing that of their commercial counterparts, including Llama-3.2-Vison and Molmo-72B.
+Additionally, we provide more detailed failure case analysis in Appendix A.6.
 Table 3: Effectiveness of Image-DPO on General VQA benchmarks.
@@
 (Bitton-Guetta et al., 2023a), and HallusionBench (Guan et al., 2023), we conduct a comparative analysis of both their high-level design principles and low-level data formats.
-This comparison incorporates qualitative and quantitative insights, offering a comprehensive understanding of their differences.
-Due to space constraints, detailed analyses are provided in Appendix A.3.
+This comparison incorporates qualitative and quantitative insights as detailed in Appendix A.4.
+Table 4: Comparisons of Image-DPO on ViLP.
+Model ViLP F ViLP P
+Score Prior Score Prior
+LLaVA-1.5-7B 29.67 71.33 37.67 65.67
++HADPO (Zhao et al., 2023b) 33.00 74.33 38.50 65.00
++RLHF-V (Yu et al., 2023a) 29.50 75.00 36.33 65.33
++EOS (Yue et al., 2024) 31.33 67.00 38.67 65.67
++SIMA (Wang et al., 2024b) 27.83 68.67 36.17 66.00
++V-DPO (Xie et al., 2024) 29.5 0 72.67 37.83 67.67
++Text-DPO 31.34 71.67 37.83 65.67
++Image-DPO 34.17 75.00 39.33 68.00
 5.1 Image DPO
@@
 The results, presented in Table 3, show consistent performance improvements across both datasets and models, further demonstrating the effectiveness of our Image DPO method.
-Table 4: Comparisons of Image-DPO on ViLP.
-Model ViLP F ViLP P
-Score Prior Score Prior
-LLaVA-1.5-7B 29.67 71.33 37.67 65.67
-+HADPO (Zhao et al., 2023b) 33.00 74.33 38.50 65.00
-+RLHF-V (Yu et al., 2023a) 29.50 75.00 36.33 65.33
-+EOS (Yue et al., 2024) 31.33 67.00 38.67 65.67
-+SIMA (Wang et al., 2024b) 27.83 68.67 36.17 66.00
-+V-DPO (Xie et al., 2024) 29.5 0 72.67 37.83 67.67
-+Text-DPO 31.34 71.67 37.83 65.67
-+Image-DPO 34.17 75.00 39.33 68.00
 TextDPO on Corrupted Images.
@@
 This work was supported by the LG AI Research grant.
-We also extend our gratitude to the OpenAI Researcher Access Program for providing the credits used to access OpenAI’s APIs.
-Finally, we greatly appreciate Chris Rockwell for his contributions to our human evaluation studies.
+We also extend our gratitude to the OpenAI Researcher Access Program for providing credits used to access OpenAI’s APIs.
+Additionally, we appreciate Chris Rockwell and Jiaming Yang for their contributions to our human evaluation studies, and thank Yongyi Yang for discussions regarding the proposition proof.
 Appendix A More details and comparisons of our benchmarks
@@
 Figure 9: One prompt we used for potential QAs designs of ViLP
-A.3 Comparisons to other datasets
+Figure 10: Realistic image comparison. Each image pair shows our original benchmark data on the left and a corresponding realistic example generated by GPT-4o on the right.
+A.3 Ablation studies: realistic images
+Our benchmark data are currently generated by DALL·E-3 (Ramesh et al., 2021) and Flux, both of which produce cartoon-like, synthetic images rather than photorealistic ones.
+To assess the impact of image style, we regenerated a subset of 45 QIA pairs using GPT-4o’s latest image generation model to enhance realism, as illustrated in Figure 10.
+We then measured changes in model correctness when these more realistic images were used, with negative values indicating performance degradation.
+For definitions of these metrics, please refer to the beginning of Section 5.
+The results in Table 5 show that increased realism slightly reduces performance for the “Score” metric in most cases, while its effects on “Prior” are generally negligible.
+These ablation findings suggest that introducing more realistic images may increase the task’s difficulty, highlighting an important direction for future research.
+$ViLP^{F}_{Prior}$ $ViLP^{F}_{Score}$ $ViLP^{P}_{Prior}$ $ViLP^{P}_{Score}$
+GPT-4o 0 -2.2% 0 0
+GPT-4o-mini 0 -1.1% 0 -2.2%
+Claude-Sonnet-3.5 -2.2% -4.4% 0 -3.3%
+Claude-Opus-3 2.2% -1.1% -2.2% -1.1%
+Table 5: Impacts of Realistic Styles.
+Each value represents the change in correctness when replacing the original images with realistic ones (Realistic - Original).
+Negative values indicate a drop in performance, suggesting increased task difficulty.
+The metric definitions are provided in the beginning of Section 5.
+A.4 Comparisons to other datasets
 Figure 11: Winoground Data Example.
@@
 While these datasets are impactful, their evaluation perspectives differ from ours, covering a range from high-level design principles to low-level formats.
-A.3.1 Compare to Winoground (Thrush et al., 2022)
+A.4.1 Compare to Winoground (Thrush et al., 2022)
 Figure 12: Example of Whoops! Dataset.Whoops! also has creative images. While unlike ours, its questions are common questions without strong language priors.
@@
 In contrast, that aspect is precisely our focus.
-A.3.2 Compare to Whoops! (Bitton-Guetta et al., 2023a)
+A.4.2 Compare to Whoops! (Bitton-Guetta et al., 2023a)
 Whoops!
@@
 In contrast, our benchmark deliberately uses suggestive questions to elicit stereotype-consistent answers, reflecting our emphasis on testing language priors.
-A.3.3 Compare to HallusionBench (Guan et al., 2023)
+A.4.3 Compare to HallusionBench (Guan et al., 2023)
 HallusionBench has two components: Visual Dependent, which focuses on testing models’ general visual reasoning skills, and Visual Supplement, which examines a model’s visual reasoning in tandem with its parametric memory.
@@
 Figure 13 showcases typical HallusionBench examples, illustrating the limited scope of both images and QAs, whereas our benchmark includes a variety of out-of-distribution images and more suggestive QAs (see Figure 1, 6, 7).
-A.4 Human Study
+A.5 Human Study
 For human evaluation, we hired Ph.D.-level candidates to participate in testing.
@@
 After the test, we updated the synonym sets for the QIAs based on their answers.
+Figure 14: Failure cases analysis. We list six typical failure cases generated by VLMs on ViLP.
+A.6 More failure cases analysis
+This subsection provides several typical failure cases observed in the inferred results of VLMs on ViLP.
+We identified several consistent failure patterns, illustrated in Figure 14.
+For deeper analysis, we also prompted the VLMs to output their reasoning before finalizing answers, allowing us to better understand how these errors arise.
+Below, we summarize the failure patterns in the same order as they appear in the figure, from left to right and top to bottom.
+- Shape recognition can fail in VLMs, causing them to revert to priors rather than accurately interpreting visual input.
+- Models sometimes struggle to count accurately. Instead of performing an actual count, they default to relying on learned priors to estimate quantities.
+- Models may refuse to accept visual information that contradicts their learned priors, whereas humans can comprehend hypothetical scenarios.
+For instance, the model recognizes the city as Beijing but rejects the correct answer because it expects the Eiffel Tower to be in Paris.
+- Sometimes the model overly relies on memorized proverbs, resulting in predictions that align with these proverbs rather than the actual content of the input image.
+- For images with creative concepts, the model may overly rely on its learned priors.
+As illustrated, a common prior is that microscopes are used to view organisms, leading the model to answer “microscope” rather than identifying the creatively depicted stapler.
+- For images with blended features, the model may rely mostly on text input while overlooking the visual cues.
+As illustrated, the VLM heavily depends on textual input leading to saxophone as the answer.
 Appendix B Image-DPO Mathematical Details
@@
 Here $\sigma$ is a logistic function. Basically, this reward function gives score jointly considering image, question and image quality.
-RL Fine-Tunning Phrase
+RL Fine-Tuning Phrase
 During the RL phase, the learned reward function is used to provide feeback to the VLM model. Following DPO paper, the optimization is formulated as :
@@
 It is the same as our objective listed in Eq.4 of the main paper.
-$\mathcal{L}(\pi_{\theta};\pi_{\text{ref}})\leq-\mathbb{E}_{(Q,I_{w},I_{l},A)\sim S}\left[\log\sigma(\alpha\frac{\pi_{\theta}(A\mid Q,I_{w})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{w})}-\alpha\frac{\pi_{\theta}(A\mid Q,I_{l})}{\pi_{\theta_{\text{ref}}}(A\mid Q,I_{l})})\right]$ (13)
+$\mathcal{L}(\pi_{\theta};\pi_{\text{ref}})\leq-\mathbb{E}_{(Q,I_{w},I_{l},A)\sim S}\left[\log\sigma(\alpha\frac{\pi_{\theta}(A\mid Q,I_{w})}{\pi_{{\text{ref}}}(A\mid Q,I_{w})}-\alpha\frac{\pi_{\theta}(A\mid Q,I_{l})}{\pi_{{\text{ref}}}(A\mid Q,I_{l})})\right]$ (13)
 In this sense, our optimization objective Eq.4 in main paper are optimizing the upper bound of RLHF, i.e., Eq.7.
@@
 To facilitate this, we use an additional text prompt (Figure 20) after the VLM generates the initial instructions (the pink region of Figure 17).
-To provide a better understanding of our generated QIAs, we randomly sampled and listed some examples of the generated QIA data, as shown in Figures 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, and 33.
+To provide a better understanding of our generated QIAs, we randomly sampled and listed some examples of the generated QIA data, as shown in Figures 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, and 35.
 Figure 17: Randomly sampled generation examples. Our proposed data generation pipeline utilizes pretrained model to generate diverse new images from existing seed image datasets.
@@
 In brief, the data generation pipeline outlined in Appendix C.1 utilizes VLMs in conjunction with pre-trained image models to generate or modify images and create corresponding question-answer pairs.
-This process results in a collection of QIA triplets, as illustrated in the Figures 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, and 33.
+This process results in a collection of QIA triplets, as illustrated in the Figures 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, and 35.
 After generating the QIA triplets, we apply three image corruption methods—Gaussian blurring, pixelation, and semantic editing—to create good-bad QIA pairs for ImageDPO training (Section 4.1), while keeping the QA components unchanged.